### v2

# PET: Preference Evolution Tracking with LLM-Generated Explainable Distribution

Abstract
Understanding how user preference evolves over time is a fundamental challenge central to modern digital ecosystems, for which Large Language Models (LLMs) are an increasingly prominent and popular approach due to their ability to comprehend the rich semantic context within behavioral data. A common practice is to use LLMs to predict a user’s next action by directly generating a ranked list of preferred items. Although effective for short-term prediction, the end-to-end generation paradigm inherently limits personalization. Its opaque decision-making process obscures holistic user profiling and exacerbates popularity bias. To address these limitations, we propose Preference Evolution Tracking (PET), a framework that reframes the task as inferring a dynamic probability distribution over a stable and interpretable lattice of preference clusters. By applying logit-probing and generative classification techniques, PET infers a user’s preference as a probability distribution, enabling transparent preference learning. On public benchmarks (Yelp, MovieLens), PET improves ranking quality by up to $40\%$ in NDCG over direct generation baselines. On a large-scale, real-world dataset from a short-video platform, it excels at ranking long-tail contents, significantly outperforming a SOTA production model by $7$ times in the NDCG score. Ultimately, PET transforms the user profile model from direct preference list generation to a transparent distributional preference mapping, paving the way for more explainable, fair, and diverse personalization systems.
1 Introduction
Understanding user interests in a transparent and interpretable manner is a central challenge in building trusted and adaptive digital ecosystems. The advent of Large Language Models (LLMs) presents a transformative opportunity in this domain. With their profound ability to comprehend context and semantics within unstructured user behavioral histories – from product reviews to content consumption sequences – LLMs promise to move beyond simple pattern matching to a deeper, more holistic understanding of user intent. The ultimate goal is not merely to predict the next action of users, but to create transparent and dynamic models of user interest that can be trusted, analyzed, and fairly acted upon.
The prevailing approach for capturing user preferences involves leveraging LLMs to directly generate ranked outputs (wang2025user; wang2024llms; ngo2024recgpt; yu2024icpl; deng2025onerec), a method proven effective for short-horizon, top-$k$ recommendation tasks. While effective for short-horizon prediction and identifying popular preferences, this direct-to-ranking paradigm leaves important gaps. First, it lacks interpretability: although LLMs encode rich internal states, these models do not expose an explicit or structured representation of the user’s overall preferences – making them difficult to audit or reuse in downstream tasks. Second, the focus on a small $k$, combined with popularity bias, often leads to the over-representation of mainstream content – neglecting the long-tail interests that define a user’s niche tastes and are critical for achieving holistic, unbiased personalization.
These limitations have significant consequences for algorithmic decision-making in user-oriented applications. The lack of an explicit user model undermines system transparency, making it difficult to govern and audit for fairness and bias. Additionally, the over-reliance on popular items not only distorts model predictions but also reinforces popularity during training – suppressing the model’s ability to learn users’ nuanced, long-tail preferences. This feedback loop amplifies the Matthew Effect (i.e., the rich get richer), reinforcing echo chambers that reduce content diversity, entrench mainstream tastes, and marginalize underrepresented creators and user segments (bakshy2015exposure; wu2024survey). This motivates a paradigm shift toward transparent, distributional user modeling, leading to our central research question:
How can we infer personalized evolving preference distribution via LLM – enhancing interpretability, fairness, and diversity?
To address this problem, we propose Preference Evolution Tracking (PET), a framework that customizes LLMs into probabilistic inference engines for modeling users’ evolving preference distributions. By leveraging pre-softmax logits over a stable preference cluster lattice (e.g., categories or tags), PET captures structured relevance signals more effectively than volatile item-level spaces. This approach enhances interpretability and diversity, especially for long-tail preferences. Specifically, PET supports two scalable inference strategies: Likelihood-based Probing, which iteratively queries the model for each cluster to build a precise distribution, and the more efficient Generative Classification, which extracts all probabilities in a single forward pass. Furthermore, PET employs Hierarchical Probing to maintain tractable inference in extreme multi-label scenarios through semantic taxonomy traversal. Theoretically, the ranking according to the inferred PET distribution guarantees an optimal ranking results (e.g., with respective to NDCG/Recall) under standard assumptions. Applied to domain-aligned LLMs, PET produces portable, interpretable user profiles for long-horizon ranking, fairness auditing, and long-tail preference modeling.
To validate PET’s effectiveness, we conduct comprehensive experiments across three diverse datasets with varying cluster complexity: MovieLens, Yelp, and a large-scale dataset from a world-leading short video platform. Our results demonstrate that the distributional approach significantly outperforms direct-generation baselines and state-of-the-art (SOTA) ranking models on public benchmarks. More critically, on the real-world data of the short video platform, our framework shows a unique ability to rank users’ niche, long-tail interests, significantly outperforming a SOTA model in production. Notably, our goal is not to replace mature production models on highly popular (head) content. Rather, PET complements them by recovering user-specific long-tail preferences that are typically underserved by popularity-optimized systems. Across all experiments, our methods prove more adept at capturing stable, long-term preferences over transient, short-term interests. Our primary contributions are threefold:
  1. 1.
    A New Paradigm for User Modeling: We introduce a novel framework that shifts the focus from optimizing short-term item predictions to generating interpretable, dynamic probability distributions over a stable lattice of human-understandable preference clusters.
  2. 2.
    A Portable and Dynamic User Profile: The learned preference distribution serves as a rich, multi-faceted user profile that captures long-tail interests and is readily portable for diverse downstream tasks, such as user segmentation, trend analysis, or fairness audits.
  3. 3.
    Theoretical and Empirical Validation: We prove that PET’s distributional ranking is theoretically superior to direct generation under standard assumptions. Empirically, we validate PET on three datasets (MovieLens, Yelp, short-video platform), achieving $40\%$ NDCG over direct generation and $20\%$ NDCG over a SOTA ranking model (Qwen3-Reranker-8B (qwen3embedding)) on public benchmarks, and a $7\times$ improvement in long-tail ranking against a SOTA production-level ranking baseline on a real-world dataset.
Figure 1: The PET framework pipeline. Left (Training): An LLM is trained on user history to learn preference distributions. Center (Inference): probing methods extracts the predicted preference distribution from the model’s internal logits. Right (Application): The transparent distribution is used for downstream tasks: personalized ranking, long-tail discovery, and interpretable user profiling.
2 Related Work
User Preference Modeling. To deliver effective personalization, user models aims to capture preferences that are both diverse and dynamic, through three key lines of work. Multi-interest models represent users via multiple embeddings linked to distinct interests (shi2023everyone; cen2020controllable; zhou2018deep). Sequential models, such as FPMC (rendle2010factorizing), GRU4Rec (hidasi2015session), SASRec (kang2018self), and BERT4Rec (sun2019bert4rec), focus on temporal dynamics to predict the next item in a user’s history. More recently, distributional modeling has become central in alignment and fairness, with methods like GDPO for group fairness (yao2024no) and RLHF-based techniques like DPL and AOT for reward uncertainty (siththaranjan2023distributional; melnyk2024distributional). While powerful, these alignment-based methods introduce significant practical considerations, as they often require costly preference-labeled datasets and complex, computationally intensive pipelines.
LLMs in Preference Modeling. Recent work has increasingly applied Large Language Models (LLMs) to the complex task of user preference modeling. Early efforts leveraged LLMs as powerful components for embedding generation (U-BERT; qiu2021u) or unified the field under a prompt-based, text-to-text paradigm (P5; geng2022recommendation). This led to end-to-end generative models like OneRec (deng2025onerec) and (ngo2024recgpt) that directly produce ranked outputs or the most relevant items (wang2025user; wang2024llms), often incorporating sophisticated alignment modules. In a parallel effort toward explainability, RecGPT uses the LLM to generate descriptive, textual user profiles (yi2025recgpt). Crucially, the direct generation of ranked lists via LLM – the approach that forms the basis of our primary baseline – has been shown to consistently outperform traditional sequential models (like SASRec (kang2018self) and BERT4Rec (sun2019bert4rec)) in various settings (geng2022recommendation; wu2024survey; zhao2024recommender; deng2025onerec).
Recent approaches also tackle complex scenarios through structured prompting and multi-context modeling, such as LLM4MSR (wang2024llm4msr), HUM (bao2025heterogeneous), and user interest exploration via latent clusters (wang2024llms). While effective for short-horizon prediction, these methods typically encode a user’s state implicitly within the model’s parameters, without surfacing an interpretable or temporally grounded preference distribution. Our work is also related to logit-probing and generative classification techniques for extracting model beliefs from pre-softmax logits (petroni2019language; schick2020s; raffel2020exploring; zhou2024quest), adapting them to infer evolving user preference distributions from fine-tuned LLMs.
3 Theoretical and Methodological Framework
In this section, we formalize the problem of inferring dynamic, comprehensive user preferences and introduce the Preference Evolution Tracking (PET) framework. We start by outlining the fundamental components of the proposed framework (see Figure 1).
3.1 Problem Formulation
We consider a setting in which the user’s interaction history up to time $t$ is represented as a sequence $H_{1:t}$. The user’s preference at any discrete time $t$ is defined as a probability distribution $\theta_{t}$ over a fixed, finite set of $K$ preference clusters $C:=\{c_{1},\cdots,c_{K}\}$:
$\theta_{t}:=(\theta_{t}(1),\cdots,\theta_{t}(K)),\text{ with }\theta_{t}(i)\geq 0\;\forall i\in[K]\text{ and }\sum_{i\in[K]}\theta_{t}(i)=1.$ (1)
We adopt a latent utility model (mcfadden1972conditional; luce1959individual; train2009discrete), a standard formulation in choice modeling, to describe how this true distribution arises. We assume that there exists a vector of unobserved “attractiveness” scores $\bm{q}:=(q_{1},\cdots,q_{K})$, where $q_{i}$ represents the true strength of the user’s preference for cluster $c_{i}$. The true preference distribution $\theta_{t}$ is then generated via a softmax function:
$\theta_{t}(i)=\frac{\exp(q_{i})}{\sum_{j\in[K]}\exp(q_{j})},\forall i\in[K].$ (2)
Our primary goal is to learn a function, parameterized by an LLM $\mathcal{M}$, that maps this interaction history to a prediction of the user’s future preference distribution, $\hat{\theta}_{t+1:T}$, for a given prediction window $\{t+1:T\}$. We choose to model preferences over this stable cluster space rather than the item space, as it is more robust to the high volatility and enormous scale of items in real-world digital ecosystems.
Since this true distribution $\theta_{t+1:T}$ is not directly observable, we construct an empirical proxy for it from the user’s interactions in a time window, which we denote as $\bar{\theta}_{t+1:T}$. Formally, for each cluster $c_{i}\in C$, its probability is the normalized frequency of interactions:
[…]
To ensure reproducibility and simplicity, we use this transparent and normalized frequency of a single action instead of heuristic weighting of different interaction types, as our framework is designed to be agnostic to the specific construction of the ground-truth label. Formally, we aim to minimize the expected divergence between these two distributions:
$\min_{\mathcal{M}}\mathbb{E}_{H_{1:t}}\left[\mathcal{D}\left(\bar{\theta}_{t+1:T}\parallel\hat{\theta}_{t+1:T}\right)\right],$ (4)
where $\mathcal{D}$ is a measure of distributional similarity, such as the Jensen-Shannon Divergence (JS-Div). The ranked list of preferences used for evaluation with metrics like NDCG is then obtained by sorting the elements of $\hat{\theta}_{t+1:T}$ in descending order.
3.2 Methodology
This section outlines the PET framework for learning user preferences, which consists of two core stages: model alignment and preference inference.
The alignment stage follows a structured two-phase training approach. First, the model undergoes large-scale pre-training (PT) on domain-specific corpora to establish foundational knowledge and capture population-level patterns. Subsequently, it is refined through supervised fine-tuning (SFT) on carefully curated user interaction sequences, effectively aligning its internal representations with the nuances of personalized preference prediction. The SFT data are constructed from each user’s full history by generating pairs (context, label), defined as follows: the context is an expanding historical window $H_{1:t}$, and the label is the user’s future preference distribution $\bar{\theta}_{t+1:T}$. This aligned model forms the basis for all subsequent preference inference methods.
We now present two inference methods for preference distribution estimation, which differ primarily in their querying strategy (more detailed design of the prompts and algorithms presented in Appendices A and B).
Method 1: Likelihood-based Probing. This approach decomposes a multi-class problem into a series of independent binary classifications. The LLM is used as a “prober” or “scorer.” For each cluster $c\in C$, we iteratively ask the LLM for inferences and extract the logit score of positive tokens and then apply a softmax function on the score vector to recover the inferred probability distribution (petroni2019language; schick2020s).
Method 2: Generative Classification. This method uses a single forward pass. A prompt is constructed that frames the task as a multi-choice question, with each cluster mapped to a unique token (e.g., ‘A’, ‘B’). We then read the model’s next-token logits for these specific cluster tokens and normalize them via a softmax function to obtain the final probability distribution (raffel2020exploring; chung2024scaling).
Extreme multi-label preference inference.
To scale PET to real-world applications involving thousands of fine-grained preference clusters – such as in large video platforms or e-commerce systems – we introduce a two-stage Hierarchical Probing strategy. Existing methods face clear limitations in such settings: Likelihood-based Probing exhibits linear computational complexity in the number of clusters $K$, becoming prohibitively expensive, while Generative Classification often yields unstable and semantically ungrounded logits when forced to assign probabilities to artificial class tokens.
To overcome these issues, we organize clusters into a two-level semantic taxonomy: coarse-grained L1 categories (e.g., Food & Drink, Technology) partition the full L2 label space into interpretable subgroups. This tree‑structured approach is also validated in recent extreme multi‑label classification work in other tasks (chen2023token; wan2023explore; zhou2024quest). Inference proceeds in two stages: (1) L1 Preference Scoping: estimates marginal probabilities over top-level categories; (2) Strategic L2 Exploration: selectively probes child clusters, guided by downstream goals. For example, one may prioritize top-ranked L1 branches for relevance, or intentionally explore long-tail branches for novelty. This decomposition improves both tractability and semantic alignment by guiding the LLM from general categories to relevant fine-grained clusters, enabling scalable and adaptive preference modeling. See the detailed computational complexity analysis in Appendix C.
3.3 Theoretical Foundations and Guarantees
This section provides a theoretical argument that PET’s methodology is optimal for standard ranking metrics compared to the direct generation paradigm. We begin with a key assumption about our model:
Assumption 1 (Isotonic Probing Estimator).
We assume that the vector of logit scores $S$ produced by PET’s probing methods (i.e., Algorithms 2 and 1) is order-preserving (isotonic) with respect to the true latent attractiveness scores $\bm{q}$. Formally, for any two clusters $c_{i},c_{j}$, we have $q_{i}\geq q_{j}\Rightarrow\mathbb{E}[S[i]]\geq\mathbb{E}[S[j]].$
1 formalizes the goal that a successfully trained model should learn scores that are correctly ordered according to the true underlying preferences, which is a foundation for many modern preference learning models (guo2017calibration; desai2020calibration; kadavath2022language; schick2020exploiting). With this, we can state the following:
Lemma 1 (Optimality of PET’s Ranking).
With 1, the ranking produced by PET by sorting its inferred probabilities maximizes the expected score for common order-aware metrics, including NDCG@k, Recall@k, and Precision@k. (See proof at Appendix D.)
Remark 1 (Sub-optimality of Direct Generation).
The sub-optimality of direct generation stems from its autoregressive nature, where the ranking is an emergent property of a sequence generation task rather than a direct readout of preference scores. This process is susceptible to multiple known issues: exposure bias, where the model’s training distribution differs from its inference distribution (ranzato2015sequence); calibration errors, where output probabilities are skewed by factors like token frequency (zhao2021calibrate); and search errors from the decoding heuristics themselves, such as beam search (wiseman2016sequence).
4 Experiment
4.1 Experiment Design and Setup
Our experimental design is crafted to comprehensively evaluate the efficacy of PET’s ability in discerning and predicting user preferences across varying granularities and long-term and short-term dynamics. We focus on using two primary inference methodologies for preference prediction and benchmark them against established and SOTA baselines.
Datasets. We evaluate PET on three real-world datasets of varying scale and complexity. MovieLens (harper2015movielens): a widely-used dataset for movie recommendations, representing a relatively small and well-defined genre (cluster) space with 19 clusters (i.e., movie genres). In total, we used 200k users with 33 million rating histories associated with 87k movies. Yelp: a large-scale dataset with 2 million users and 8 million ratings across 1,311 business categories, providing a significantly higher-dimensional preference space. A Real-World Short-Video Dataset: this proprietary dataset is derived from the a real-world short video platform and captures large-scale real-world user interaction behaviors in a highly dynamic environment, reflecting diverse and rich user preference patterns. To ensure privacy, all identifiers are rigorously anonymized. The dataset comprises 50 million implicit feedback records from 16k randomly sampled users, covering 15 million items. Each record contains only anonymized user IDs, item IDs, and interaction types.
[…]
Models and Methods. We use two families of open-source 8B-parameter language models: Qwen3-8B (qwen3technicalreport) and DeepSeek-Distill-Llama-8B (deepseekai2025deepseekr1incentivizingreasoningcapability), selected for their strong balance of SOTA performance and computational efficiency. Each model is trained with our PT+SFT (LoRA fine-tune (hu2022lora)) pipeline. This setup supports a comprehensive ablation of training strategies. Our experiments evaluate the primary inference methodologies detailed in Section 3: Likelihood-based Probing (Algorithm 1) and Generative Classification (Algorithm 2). For the Yelp dataset with large cluster space, we specifically evaluate our Hierarchical Probing approach (Algorithm 3). For all methods, the final inferred probability distribution is used to generate a ranked list of clusters for evaluation.
Baselines. We compare PET with two strong LLM-based baselines. To isolate the effect of our inference paradigm, the first baseline Direct Generation uses the same pre-trained and fine-tuned LLM as PET but is prompted to directly generate a top-k ranked list of preferences without producing a probability distribution (geng2022recommendation; deng2025onerec; wang2025user; wang2024llms). As this method does not produce a probability distribution, it is evaluated only on ranking metrics. The second baseline is Qwen3-Reranker-8B, a SOTA pre-trained model specifically optimized for ranking tasks (qwen3embedding). We use it out of the box without any additional fine-tuning. Given a user history and a candidate cluster space, it produces both a ranked list and a probability distribution. For the short-video dataset, we compare against the platform’s production Transformer-based ranker model.
Evaluation Metrics. Model performance is comprehensively evaluated using established metrics compared to long-term (LT) and short-term (ST) ground truths, with a focus on ranking quality and distributional similarity. Ranking metrics include NDCG@k, which assesses the quality of ranking order with an emphasis on top positions; Recall@k, measuring the proportion of relevant items captured within the top k results; and Precision@k, evaluating the exactness of top-k recommendations. Distribution similarity is quantified using Jensen-Shannon Divergence (JS-Div), where lower values indicate closer alignment between predicted and actual preference distributions. All results are averaged across test samples and reported per context window size as specified in the dataset.
Figure 2: Comprehensive comparison of different methods on MovieLens dataset across a range of context windows (i.e., 1,3,5, and 8 sessions). Prediction windows: long-term and short-term. Metrics: NDCG@$1,5,10$ and JS-Divergence. We compare our Logit Probing and Generative Classification methods against (1) Direct Generation benchmark trained (PT+SFT) on a Qwen3-8B base model and (2) a SOTA Qwen3-Reranker-8B ranking model. Note that for the Direct Generation baseline, the JS-divergence will not be available. Sample size: 38,434.
4.2 How Effective and Robust is PET across Different Context Horizon?
We begin by evaluating PET on the MovieLens dataset with 19 clusters, aiming to systematically test its effectiveness and robustness across following key dimensions: (1) different input history lengths and (2) prediction horizons (short-term vs. long-term). We vary the input context window from 1 to 8 sessions, and assess performance on both short-term (next-session) and long-term (multi-session) preference prediction.
As shown in Figure 2, the results demonstrate the robustness and effectiveness of PET across a wide range of settings. The performance gains are especially pronounced in full-list metrics (NDCG@5, NDCG@10), where PET’s distributional inference enables richer modeling of user preferences beyond top-1 accuracy. For example, with Qwen3-8B, context window =8=8, and long-term prediction, Logit-Probing achieves NDCG@10 = 0.863, with a +55% improvement over Direct Generation and +20% gain over Qwen3-Reranker-8B. It also improves distributional alignment, reducing JS-Div to 0.316, a 22% relative reduction compared to the Qwen3-Reranker-8B’s 0.406. While Direct Generation is occasionally competitive at NDCG@1, its performance deteriorates as the number of predicted clusters increases, as it might suffers from exposure bias, noted in Remark 1. Our Generative Classification method also shows consistent improvements over Direct Generation on Qwen3-8B base model. Though slightly less accurate than Logit-Probing, it offers a compelling latency–accuracy trade-off due to its single-pass nature, making it a strong candidate for real-time or resource-constrained environments. (See more experimental results in other settings in Appendix F)
To demonstrate PET’s ability to build interpretable profiles, we sampled 125 users and divided their histories into four equal time periods. Using PET’s Likelihood-based Probing, we inferred a distribution over 19 genres for each user in each period. Aggregating these across users yielded group-level preference distributions that evolve over time – visualized in both a heatmap and 3D ribbon plot (Figure 3). These results show that user preference distributions are not static but migrate in nuanced ways. Dominant genres like Drama, Comedy, and Action exhibit stability across periods, representing structural, long-term user interests. In contrast, genres such as Adventure, Romance, and Sci-Fi demonstrate significant temporal fluctuations – highlighting PET’s ability to capture volatile, context-dependent preferences. Notably, genres like Sci-Fi, Fantasy, and IMAX show gradual upward shifts, possibly reflecting emerging or re-surfacing interests. Together, these dynamics validate PET’s unique capability not only in preference prediction, but also in generating descriptive, evolving group-level user profiles – an essential feature for diagnosing user behavior shifts, designing cold-start interventions, or understanding preference stability versus exploration.
4.3 Can PET Handle Thousands of Clusters?
To evaluate PET’s scalability in high-dimensional settings, we use the Yelp dataset with 1,311 fine-grained categories. Since exhaustive Logit-Probing becomes intractable at this scale, we adopt a two-stage Hierarchical Probing strategy (Algorithm 3). First, we construct a semantic hierarchy that maps all categories into 26 high-level L1 clusters (e.g., “Food & Restaurants,” “Health & Medical”). For demonstration, we focus on the high-traffic L1 cluster “Food & Restaurants,” further decomposed into 19 L2 sub-clusters (e.g., “Mexican,” “Vegan”). During inference, PET first identifies a user’s coarse L1 interest, then performs fine-grained ranking within the selected L2 cluster. While operators may target different L1 clusters depending on downstream needs, this experiment provides one representative use case. (See detailed mapping construction in Appendix E.)
Table 1 shows that PET with Hierarchical Probing achieves strong performance across both coarse (L1) and fine-grained (L2) levels. At L1, Logit-Probing is unequivocally dominant. It achieves high and robust NDCG scores (e.g., NDCG@1,5,10 are 0.980, 0.879, and 0.915 for Qwen3-8B), demonstrating an exceptional ability to identify a user’s general interests. In stark contrast, both the Direct Generation and Generative Classification methods fail at this stage, with performance often near zero. These results reinforce the necessity of iterative probing in high-dimensional settings where broad disambiguation is critical. At L2, where the ranking task becomes more localized, both Logit-Probing and Generative Classification significantly outperform the Direct Generation benchmark. Notably, while Logit-Probing remains the top performer for Qwen3-8B (0.666 NDCG@10), Generative Classification emerges as the clear winner for the DeepSeek-distill-Llama-8B model (0.574 NDCG@10), suggesting that decoding strategies for fine-grained inference can be model-dependent.
Together, these findings demonstrate that PET, equipped with hierarchical probing, scales effectively to large cluster spaces. The coarse-to-fine process not only preserves tractability but also allows flexible adaptation across model architectures — with Logit-Probing excelling in breadth and Generative Classification offering a strong, efficient alternative at finer levels of granularity.
(a) Group-level preference evolution as a probability heatmap
(b) Group-level preference evolution as a bar chart
Figure 3: Evolution of group-level movie genre preferences as probability distribution across four time periods. Dataset: Movielens, 125 users over 4 periods. Method: PET Likelihood-based Probing with PT+SFT on Qwen3-8B. Note: due to limited space, we only show top-10 clusters (genres) here.
Qwen3-8B DeepSeek-distill-Llama-8B
Level Method NDCG@1 NDCG@5 NDCG@10 NDCG@1 NDCG@5 NDCG@10
L1 (26 clusters) Direct Generation .008 .139 .330 .016 .308 .343
Logit Probing .980 .879 .915 .925 .856 .880
Generative Classification .107 .118 .163 .001 .062 .119
L2 (19 sub-clusters) Direct Generation .298 .368 .509 .200 .153 .329
Logit Probing .513 .562 .666 .261 .268 .363
Generative Classification .466 .349 .454 .684 .553 .574
Table 1: Hierarchical Probing’s performance on the high-dimensional Yelp dataset for the long-term prediction task. Results are shown for a context window of 8 using the PT+SFT pipeline. Best NDCG scores for each level are in bold. L1 category contains 26 clusters; L2 category contains 19 clusters under one L1 category of “Food & Restaurants”. Alternative taxonomies can be easily integrated depending on platform objectives. Sample Size: 5000.
4.4 How Does PET Perform on Real-World Short-Video Platform?
We conclude our empirical study with a large-scale, private short-video dataset from a major content platform, comprising 78 content clusters. This experiment is designed to test our method’s efficacy in a real-world environment against a heavily optimized, production-level Transformer-based ranking model. The primary challenge in such an ecosystem is overcoming the inherent feedback loop, where users primarily interact with content they are already shown. Therefore, our evaluation focuses on PET’s ability to complement an existing production model that is heavily optimized for popular, high-frequency head content. In contrast, PET is designed to surface niche, long-tail interests that are often suppressed by popularity bias. We aggregate interactions into daily sessions (a 24-hour window), ensuring a denser and more comprehensive view of a user’s preferences. All results use the PT+SFT pipeline with Qwen3-8B. Following industry practice, we use play duration – the total time a user spends watching each content item – as a robust and continuous proxy for engagement. (See more details in Section E.2)
Results in Table 2 showcase PET’s effectiveness in promoting long-tail interest, diversity, and personalized recommendation. We highlight the setting of 30-day history to 14-day prediction, as it represents the most challenging long-range forecasting task. The production model fails to identify users’ long-tail preferences, with an NDCG@20 of just 0.0243 on the long-tail segment. PET with Logit-Probing achieves NDCG@20 of 0.1971 on the same tail – a +711% improvement. This demonstrates PET’s ability to surface underrepresented, user-specific content that is typically overlooked by popularity-driven production models. Notably, results are consistent for $k\in\{1,5,10,20\}$ with PET achieving more gains compared to the production model when $k$ is small. We also compare across prediction horizons and observe that PET slightly outperforms the baseline more significantly when the forecasting window is longer (14 days vs. 7 days), consistent with our earlier findings at Section 4.2 regarding LLMs’ strength in modeling long-term user preferences. Finally, PET’s performance remains robust across different context lengths. It maintains stable ranking quality even with limited user history, underscoring its ability to effectively summarize user interests with sparse or noisy data.
Taken together, these findings validate PET as a practical and scalable complement to production systems. enhancing long-tail coverage and personal relevance without sacrificing overall performance. Its ability to enhance diversity and personalization – especially for the most engaged users – marks a valuable addition to real-world recommender systems.
Context Prediction Method NDCG@1 NDCG@5 NDCG@10 NDCG@20
14 Days 7 Days SOTA Production (Production) .007 .009 .012 .034
Logit-Probing .113 .124 .147 .203
14 Days SOTA Transformer (Production) .004 .006 .008 .024
Logit-Probing .110 .126 .150 .208
30 Days 7 Days SOTA Transformer (Production) .007 .009 .012 .034
Logit-Probing .103 .118 .138 .192
14 Days SOTA Transformer (Production) .004 .006 .008 .024
Logit-Probing .108 .122 .143 .197
Table 2: Long-tail preference learning performance on the short-video dataset for high-activity users, comparing PET’s Logit Probing against a SOTA production model across various history (i.e., 14 and 30 days) and prediction windows (i.e., 7 and 14 days). All results are from the PT+SFT pipeline with the Qwen3-8B model.
5 Conclusion
In this paper, we introduce Preference Evolution Tracking (PET), a framework that casts LLM-based personalization as distributional preference mapping: PET infers a user’s cluster-level probability vector via Likelihood-based Probing and Generative Classification, and scales to large label spaces through Hierarchical Probing. Under a mild isotonicity assumption, ranking by PET’s inferred probabilities is optimal for standard order-aware metrics. Empirically, PET is robust across datasets and horizons, outperforming strong generative and reranking baselines and substantially improving coverage of long-tail interests on an industrial-scale corpus. Beyond accuracy, PET yields interpretable, temporally evolving user profiles that enable group-level user analysis and preference-shift detection. Future work includes extending PET to model richer user representations by incorporating auxiliary signals such as demographic attributes, behavioral embeddings, and multimodal content. These enhancements aim to enable more personalized, adaptive, and interpretable preference modeling in real-world recommendation systems.
Appendix
Appendix A Algorithms
Algorithm 1 Likelihood-based Probing.
1:
[…]
9:Predicted preferences $\hat{\theta}_{t+1:T}$ over a future period.
[…]
11:for $j\leftarrow 1$ to $K$ do
12:  Use $T_{\text{probe}},H_{1:t},c_{j}$ to construct a prompt $p_{j}$. $\triangleright$ Create prompt for cluster $c_{j}$
13:  $\mathbf{z}_{j}\leftarrow\mathcal{M}(p_{j})$ $\triangleright$ Get logit vector $\mathbf{z}_{j}\in\mathbb{R}^{|V|}$ for prompt $p_{j}$
14:  $s_{+}\leftarrow\sum_{v\in V_{+}}(\mathbf{z}_{j}[v])/|V_{+}|$ $\triangleright$ Calculate mean logit score for all affirmative tokens
15:  $s_{-}\leftarrow\sum_{v\in V_{-}}(\mathbf{z}_{j}[v])/|V_{-}|$ $\triangleright$ Calculate mean logit score for all negative tokens
16:  $S[j]\leftarrow\frac{\exp(s_{+})}{\exp(s_{+})+\exp(s_{-})}$ $\triangleright$ Compute ’yes’ probability as the cluster’s score
17:end for
[…]
Algorithm 2 Generative Classification.
1:
[…]
8:Predicted preferences $\hat{\theta}_{t+1:T}$ over a future period.
[…]
10:Use $T_{\text{gen}},H_{1:t},C$ to construct a prompt pp.
[…]
12:for $j\leftarrow 1$ to $K$ do
13:  $S[j]\leftarrow\mathbf{z}[v_{j}]$ $\triangleright$ Extract the raw logit for the token of cluster $c_{j}$
14:end for
[…]
Algorithm 3 Hierarchical Probing with Strategic Exploration.
1:
[…]
7:Predicted preferences $\hat{\theta}$ over the L2 clusters $C_{L2}$.
8:
[…]
10:Use $H_{1:t}$ and $C_{L1}$ to get L1 scores $S_{L1}$ via “Likelihood-based Probing” (Algorithm 1).
[…]
12:
[…]
16:for each selected cluster $c_{j,L1}$ in $C_{L1,\text{selected}}$ do
17:  $C_{L2,\text{subset}}\leftarrow\text{map}(c_{j,L1})$ $\triangleright$ Get L2 children for this L1 parent.
18:  Construct a conditional prompt for $C_{L2,\text{subset}}$ (e.g., “Given interest in $c_{j,L1}$, …”).
19:  Use the conditional prompt to get scores $S_{L2,\text{subset}}$ via “Likelihood-based Probing”.
20:  $P_{L2|L1}\leftarrow\text{softmax}(S_{L2,\text{subset}})$ $\triangleright$ Get conditional probabilities $P(C_{L2}|c_{j,L1})$.
21:  for each cluster $c_{k}$ in $C_{L2,\text{subset}}$ do
22:   $S_{L2}[k]\leftarrow P_{L2|L1}[c_{k}]\times P_{L1}[j]$ $\triangleright$ Apply chain rule for joint probability.
23:  end for
24:end for
25:
[…]
Appendix B Prompt of Inference
Prompt for Likelihood-based Probing Algorithm 1
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Considering the user’s long-term preferences from their movie rating history, do they like {GENRE} movies? Answer in “Yes” or “No”.
Prompt for Generative Classification Algorithm 2
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Considering the user’s long-term preferences from their movie rating history, which genre do they like MOST? Answer with the letter only (A, B, C, etc.):
Prompt for Direct Generation (top-$1$)
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Question: Based on the user’s long-term preferences from their entire history, tell me the cluster they like the MOST. Answer with the letter only (A, B, C, etc.):
Prompt for Direct Generation (top-$3$)
User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
Time 2: rated “The Godfather” 5/5 (Crime, Drama);
Time 3: rated “Toy Story” 4/5 (Animation, Comedy)
Question: Based on the user’s long-term preferences from their entire history, rank the top 3 genres they like the most. Answer with the letter only (A, B, C, etc.):
Appendix C Computational Complexity Analysis
Notation.
Let $K$ be the number of preference clusters and $T$ the prompt length (tokens). Let $C(\mathcal{M},T)$ denote the cost of a single forward pass of the language model $\mathcal{M}$ on a prompt of length $T$ (the dominant compute). Each cluster $c$ is associated with a fixed “verbalizer” (token or short phrase) used to read its score from next-token logits.
Complexity of PET Methods
Likelihood-based Probing Algorithm 1.
Our implementation is iterative: we issue a focused prompt per cluster, effectively scoring clusters independently. The total cost scales linearly with the number of probes:
$\text{Cost}\;\approx\;O\!\big(K\cdot C(\mathcal{M},T)\big).$
Generative Classification Algorithm 1.
This method is single-shot: one prompt, one forward pass, then read the next-token logits for the $K$ verbalizers and normalize. The cost is constant with respect to $K$:
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)\big).$
Hierarchical Probing Algorithm 3.
This two-stage method targets large $K$. Let $K_{1}$ be the number of coarse L1 clusters, $B$ the number of L1 branches explored, and $K_{2,\text{avg}}$ the average number of L2 sub-clusters per explored branch. Using iterative probing at both levels, the total cost is the sum of L1 scoping and L2 exploration:
$\text{Cost}\;\approx\;O\!\big(K_{1}\cdot C(\mathcal{M},T_{L1})\big)\;+\;O\!\big(B\cdot K_{2,\text{avg}}\cdot C(\mathcal{M},T_{L2})\big).$
Since typically $K_{1}\ll K$ and $B\cdot K_{2,\text{avg}}\ll K$, this is substantially cheaper than flat iterative probing $O\!\big(K\cdot C(\mathcal{M},T)\big)$.
Complexity of Baselines (for comparison)
Direct Generation (top-$k$ list).
One prompt with autoregressive decoding of approximately $O(k)$ output tokens:
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)+k\big).$
If re-prompted separately for multiple cutoffs (e.g., $k=1,5,10$), end-to-end latency increases linearly with the number of cutoffs.
Cross-encoder Reranker.
Score each $(\text{history},\text{cluster})$ pair independently:
$\text{Cost}\;\approx\;O\!\big(K\cdot C(\mathcal{M},T)\big).$
Practical Notes
  • KV cache reuse. For iterative methods, reusing the encoded history reduces repeated compute; most marginal cost is in the final token(s) where logits are read.
  • Batching. Batching improves throughput for iterative probing and reranking, but tail latency still scales with the number of batches/probes.
  • Multi-$k$ evaluation. PET’s distributional methods (Generative Classification; Hierarchical with small $B$) yield a full distribution in the same pass(es), so reporting multiple $k$ does not require re-decoding.
Appendix D Proofs
For Lemma 1:
Proof.
By 1, the scores $S$ from PET preserve the order of the true latent attractiveness scores $\bm{q}$. Since the softmax function is monotonic, the inferred probability distribution $\hat{\theta}$ also preserves this order. Thus, sorting by PET’s inferred probabilities produces a ranking, $\pi_{\text{PET}}$, that is identical to the optimal ranking, $\pi^{*}$, which sorts by the true latent scores $\bm{q}$. For a given user and within the time window of $\{t+1:T\}$, we let the vector of attractiveness score $\bm{q}$ be the true latent relevance score of a user for the cluster space. And we denote the sorted (in descending order) vector of $\bm{q}$ as $\tilde{\bm{q}}$ with $\tilde{q}_{1}\geq\cdots\geq\tilde{q}_{K}$. By 1, the retried logit score vector preserves the same order $S$ as $\bm{q}$. Thus, as the softmax function of $S$ preserves the same order of $S$, then by the procedure of PET – sorting the inferred probability distribution in the descending order – the outputted rank by the PET’s procedure has the exactly the same order as the $\tilde{\bm{q}}$.
On the other hand, let’s visit the definition of three ranking metrics. For any given ranking/permutation of cluster set $C$, denoted as $\pi$ (with the associated score vector $q_{\pi}$), we define
  • Recall@k: $\{\text{relevant items in top}-k\}/K$.
[…]
  • NDCG@k: $\text{DCG}(\pi)@k/\text{IDCG}@k$, where $\text{DCG}(\pi)@k:=\sum_{r+1}^{K}\frac{q_{\pi}(r)}{\log_{2}(r+1)}$ and $\text{IDCG}@k$ is the best possible $\text{DCG}(\pi)@k$ over all possible permutation $\pi$.
It is trivial to see that, all these three ranking metrics are maximized when the permutation $\pi$ has the same order as $\tilde{\bm{q}}$. Since the outputted rank by PET has the same order as $\tilde{\bm{q}}$, then PET achieves the optimal ranking scores with respect to these three ranking methods, which gives equal or better rank scores than the direct generation method. ∎
Appendix E Detailed Experimental Design
E.1 Yelp Dataset
On the Yelp dataset, we evaluate PET in a high-dimensional setting with 1,311 fine-grained categories in the Yelp dataset, designed to test the scalability of our approach. In this regime, Logit Probing becomes prohibitively expensive due to the large output space. To address this, we adopt Hierarchical Probing (Algorithm 3) that decomposes the task into two levels: a coarse-grained cluster identification followed by fine-grained ranking within the selected scope.
We construct a hierarchical mapping that maps 1,311 categories into 26 high-level L1 clusters (e.g., “Food & Restaurants”, “Health & Medical”). For a high-traffic category like “Food & Restaurants,” we then define a corresponding L2 space with 19 sub-clusters (e.g., “Mexican,” “Vegan”) under L1’s “Food & Restaurants”. In inference, PET first identifies the user’s broad L1 interests and then “zooms in” to perform a nuanced L2 ranking within the top-predicted L1 category. While operators may freely select which high-level categories to probe based on application needs, this experiment offers one representative setup on the Yelp dataset.
Hierarchical Probing Mapping. We ask Gemini-2.5 Pro and GPT-5 to produce this mapping with the prompt:
You are an expert in ontology construction, taxonomy design, and user preference modeling, with extensive experience in building interpretable, semantically meaningful cluster hierarchies for preference learning. I will provide you a list of raw user preference clusters, each with an associated weight or frequency representing its importance or user interaction volume (e.g., number of views). Bicycles: number of reviews xxx, average ratings xxx;
Massage: number of reviews xxx, average ratings xxx;
Your task is to: 1. Group these raw clusters into high-level categories (Level-1) that are semantically meaningful and interpretable. 2. Assign each raw cluster to one and only one Level-1 category, forming a two-level hierarchy. 3. Return the result as a JSON object, where the keys are Level-1 category names and the values are lists of Level-2 clusters (i.e., raw cluster names). You should aim for: 1. 20 to 26 Level-1 categories total. 2. Coherent semantics within each group. 3. Human-readable, concise names for each Level-1 category (e.g., “Food & Drink”, “Fitness”, “Beauty”).
With the mapping, of which we verified that Level-1 categories are well-balanced, with no single group dominating the total frequency or collapsing into overly broad labels:
L1 Categories: Active Life & Fitness, Active Life, Sports & Recreation, Arts, Entertainment & Events, Automotive, Beauty & Spas, Cannabis Services, Community & Government, Education, Event Planning & Services, Farms & Ranches, Food & Restaurants, Health & Medical, Home & Public Services, Home Maintenance, Hotels & Travel, Internet & Communications, Local & Public Services, Nightlife & Bars, Personal Services, Pets, Professional & Financial Services, Real Estate, Religious & Community, Shopping & Retail, Specialty Shops, Specialty Vehicles L2 Categories: Bicycles, Bike Shop, Active Life, Aerial Fitness, Airsoft, Amateur Sports Teams, Archery, Badminton, Barre Classes, Baseball Fields, Basketball Courts, Beach Bars, Beaches, Bicycle Paths, Bike Parking, Bike Sharing, Bike tours, Bikes, Bocce Ball, Boot Camps, Boxing, Brazilian Jiu-jitsu, …
E.2 Real-World Short-Video Dataset
On the short-video dataset, we select the play duration as the proxy metric to represent the user interests. Compared to sparse binary signals like likes or saves, play duration offers a more reliable and verifiable measure of user interest. While some approaches construct weighted combinations of multiple engagement signals, we deliberately opt for a single, interpretable metric to avoid introducing volatility and platform-specific heuristics. The ground-truth distribution is defined as the aggregated play duration across future windows (3, 7, or 14 days). Although we cannot release raw correlation statistics due to data sensitivity and privacy constraints, internal audits confirm strong alignment between play duration and user feedback.
Our analysis targets high-activity users ($\geq$21 active days in the past month), as they represent the platform’s core, most engaged cohort. Since we only have $30$-days of user data, we pick high-activity users for their dense acitivities.
Durations are normalized per user to form a probability distribution over clusters, preventing heavy-usage magnitude from confounding preference shares.
Appendix F Additional Experimental Results
F.1 Evaluation of PET on Movielens
Inference procedure of Direct Generation. Unlike PET, which infers a full probability distribution over all 19 clusters, Direct Generation is prompted separately for each top-k values.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method NDCG@1 NDCG@5 NDCG@10 JS Div NDCG@1 NDCG@5 NDCG@10 JS Div
1 Long A: Direct Generation .813 .573 .644 .609 .440 .540
B: Logit-Probing .781 .776 .823 .320 .734 .786 .840 .373
C: Gen. Class. .795 .725 .736 .547 .562 .606 .632 .462
Qwen3-Reranker-8B .814 .713 .728 .399
Short A: Direct Generation .701 .506 .571 .499 .369 .450
B: Logit-Probing .667 .675 .740 .466 .542 .619 .714 .514
C: Gen. Class. .638 .661 .789 .538 .499 .490 .501 .551
Qwen3-Reranker-8B .650 .566 .652 .525
3 Long A: Direct Generation .820 .573 .644 .601 .428 .533
B: Logit-Probing .815 .815 .858 .309 .758 .812 .812 .370
C: Gen. Class. .830 .760 .758 .547 .637 .624 .624 .462
Qwen3-Reranker-8B .809 .714 .727 .401
Short A: Direct Generation .702 .495 .554 .492 .361 .453
B: Logit-Probing .686 .674 .707 .460 .501 .612 .712 .507
C: Gen. Class. .661 .631 .671 .539 .492 .499 .507 .545
Qwen3-Reranker-8B .624 .563 .645 .525
5 Long A: Direct Generation .840 .538 .629 .604 .428 .542
B: Logit-Probing .830 .815 .825 .315 .763 .817 .867 .367
C: Gen. Class. .847 .777 .764 .547 .423 .528 .551 .463
Qwen3-Reranker-8B .802 .709 .725 .403
Short A: Direct Generation .723 .496 .546 .493 .359 .454
B: Logit-Probing .716 .720 .781 .460 .496 .606 .708 .507
C: Gen. Class. .673 .635 .676 .535 .489 .498 .507 .546
Qwen3-Reranker-8B .606 .552 .641 .525
8 Long A: Direct Generation .848 .530 .613 .605 .425 .546
B: Logit-Probing .831 .822 .863 .316 .774 .823 .872 .367
C: Gen. Class. .846 .760 .762 .542 .362 .518 .542 .465
Qwen3-Reranker-8B .801 .701 .721 .406
Short A: Direct Generation ..711 .499 .529 .492 .359 .456
B: Logit-Probing .710 .719 .780 .462 .490 .609 .711 .507
C: Gen. Class. .657 .632 .673 .530 .486 .497 .509 .547
Qwen3-Reranker-8B .609 .522 .638 .528
Table 3: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A) and Qwen3-Reranker-8B. Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Precision@1 Precision@5 Precision@10 Precision@1 Precision@5 Precision@10
1 Long A: Direct Generation .996 .847 .779 .979 .857 .803
B: Logit-Probing .993 .951 .754 .993 .966 .919
C: Gen. Class. .992 .890 .736 .948 .821 .760
Short A: Direct Generation .921 .549 .484 .772 .467 .417
B: Logit-Probing .887 .711 .589 .788 .688 .608
C: Gen. Class. .882 .627 .517 .772 .588 .434
3 Long A: Direct Generation .999 .857 .773 .972 .845 .799
B: Logit-Probing .991 .960 .914 .996 .973 .926
C: Gen. Class. .997 .911 .792 .897 .854 .755
Short A: Direct Generation .926 .555 .488 .775 .466 .429
B: Logit-Probing .909 .738 .617 .755 .693 .618
C: Gen. Class. .902 .654 .518 .775 .601 .446
5 Long A: Direct Generation .998 .866 .767 .976 .842 .804
B: Logit-Probing .996 .961 .913 .993 .973 .931
C: Gen. Class. .998 .906 .789 .869 .857 .756
Short A: Direct Generation .930 .559 .488 .768 .463 .429
B: Logit-Probing .926 .749 .622 .746 .687 .620
C: Gen. Class. .912 .662 .525 .762 .597 .443
8 Long A: Direct Generation .999 .869 .776 .975 .837 .807
B: Logit-Probing .996 .963 .918 .998 .977 .934
C: Gen. Class. .999 .907 .794 .858 .861 .757
Short A: Direct Generation .925 .555 .486 .774 .461 .431
B: Logit-Probing .922 .749 .621 .736 .694 .622
C: Gen. Class. .907 .658 .521 .762 .597 .447
Table 4: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold.
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
1 Long A: Direct Generation .072 .299 .547 .071 .301 .563
B: Logit-Probing .072 .342 .637 .073 .350 .658
C: Gen. Class. .072 .317 .552 .068 .289 .533
Short A: Direct Generation .124 .349 .603 .097 .285 .519
B: Logit-Probing .120 .461 .743 .106 .446 .772
C: Gen. Class. .118 .402 .654 .097 .367 .538
3 Long A: Direct Generation .073 .304 .544 .070 .297 .562
B: Logit-Probing .072 .345 .651 .073 .354 .666
C: Gen. Class. .072 .325 .559 .064 .302 .531
Short A: Direct Generation .124 .354 .608 .096 .278 .527
B: Logit-Probing .122 .477 .777 .101 .443 .773
C: Gen. Class. .121 .417 .653 .096 .369 .543
5 Long A: Direct Generation .073 .310 .544 .070 .295 .563
B: Logit-Probing .073 .349 .656 .072 .351 .666
C: Gen. Class. .073 .326 .562 .061 .302 .529
Short A: Direct Generation .125 .357 .609 .095 .278 .529
B: Logit-Probing .125 .486 .784 .099 .441 .780
C: Gen. Class. .123 .423 .662 .095 .368 .542
8 Long A: Direct Generation .073 .309 .547 .070 .292 .564
B: Logit-Probing .073 .348 .655 .073 .353 .668
C: Gen. Class. .073 .324 .562 .060 .304 .529
Short A: Direct Generation .125 .356 .607 .096 .277 .532
B: Logit-Probing .126 .488 .785 .099 .445 .781
C: Gen. Class. .123 .422 .659 .095 .370 .547
Table 5: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold.
F.2 Evaluation of PET on Yelp
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method NDCG@1 NDCG@5 NDCG@10 JS Div NDCG@1 NDCG@5 NDCG@10 JS Div
1 Long A: Direct Generation 0.0015 0.1069 0.2430 0.0127 0.3109 0.3425
B: Logit-Probing 0.8604 0.8436 0.8669 0.6145 0.9581 0.8501 0.8777 0.6499
C: Gen. Class. 0.0945 0.1147 0.1587 0.6969 0.0000 0.1021 0.1437 0.7245
Short A: Direct Generation 0.0149 0.1133 0.2686 0.0127 0.2971 0.3333
B: Logit-Probing 0.7430 0.7467 0.7891 0.6768 0.7072 0.7352 0.7650 0.6969
C: Gen. Class. 0.0398 0.1122 0.1549 0.7301 0.0001 0.0928 0.1468 0.7296
3 Long A: Direct Generation 0.0040 0.1164 0.3381 0.0170 0.3142 0.3450
B: Logit-Probing 0.9557 0.8724 0.9016 0.5976 0.9590 0.8683 0.8866 0.6527
C: Gen. Class. 0.1254 0.1230 0.1684 0.6935 0.0000 0.0647 0.1451 0.7179
Short A: Direct Generation 0.0277 0.1403 0.3556 0.0219 0.3034 0.3379
B: Logit-Probing 0.8743 0.7970 0.8548 0.6611 0.3391 0.4507 0.5194 0.7020
C: Gen. Class. 0.0454 0.1120 0.1577 0.7235 0.0028 0.0552 0.1823 0.7334
5 Long A: Direct Generation 0.0115 0.1279 0.3522 0.0176 0.3136 0.3481
B: Logit-Probing 0.9629 0.8713 0.9075 0.5848 0.7814 0.8057 0.8350 0.6579
C: Gen. Class. 0.1204 0.1243 0.1720 0.6839 0.0004 0.0567 0.1554 0.7143
Short A: Direct Generation 0.1101 0.1472 0.3557 0.0203 0.3001 0.3394
B: Logit-Probing 0.8823 0.7923 0.8585 0.6511 0.6402 0.7237 0.7643 0.6944
C: Gen. Class. 0.0411 0.1137 0.1556 0.7194 0.0905 0.0802 0.1107 0.7317
8 Long A: Direct Generation 0.0081 0.1390 0.3297 0.0156 0.3084 0.3429
B: Logit-Probing 0.9795 0.8787 0.9152 0.5734 0.9246 0.8563 0.8796 0.6546
C: Gen. Class. 0.1070 0.1183 0.1628 0.7012 0.0013 0.0615 0.1187 0.7168
Short A: Direct Generation 0.0739 0.1347 0.3152 0.0176 0.2948 0.3367
B: Logit-Probing 0.8593 0.7874 0.8599 0.6470 0.6809 0.7250 0.7637 0.6984
C: Gen. Class. 0.0402 0.1227 0.1526 0.7341 0.1095 0.0863 0.1292 0.7328
Table 6: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: NDCG. Best results for each setting are in bold. Level: L1
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Precision@1 Precision@5 Precision@10 Precision@1 Precision@5 Precision@10
1 Long A: Direct Generation 0.0054 0.4100 0.2583 0.1423 0.2503 0.2667
B: Logit-Probing 0.9538 0.6749 0.4897 0.9892 0.5025 0.3898
C: Gen. Class. 0.6152 0.3920 0.3456 0.0018 0.3508 0.2656
Short A: Direct Generation 0.0408 0.2252 0.1707 0.0444 0.2130 0.1885
B: Logit-Probing 0.8613 0.3718 0.2754 0.7905 0.3910 0.2628
C: Gen. Class. 0.1453 0.2006 0.1947 0.0006 0.1976 0.1681
3 Long A: Direct Generation 0.0186 0.3880 0.2929 0.1531 0.2516 0.2672
B: Logit-Probing 0.9886 0.6379 0.4957 0.9946 0.5557 0.4017
C: Gen. Class. 0.7713 0.3863 0.3482 0.0006 0.2281 0.2606
Short A: Direct Generation 0.0828 0.2101 0.2060 0.0576 0.2172 0.1902
B: Logit-Probing 0.9526 0.3905 0.3186 0.4106 0.2403 0.2018
C: Gen. Class. 0.1477 0.1963 0.1983 0144 0.1005 0.1634
5 Long A: Direct Generation 0.0690 0.3539 0.2918 0.1609 0.2549 0.2881
B: Logit-Probing 0.9898 0.6293 0.5030 0.9040 0.5651 0.4384
C: Gen. Class. 0.7863 0.4104 0.3592 0.0018 0.1878 0.1997
Short A: Direct Generation 0.3854 0.1894 0.2015 0.0540 0.2172 0.1975
B: Logit-Probing 0.9616 0.3864 0.3268 0.7419 0.4715 0.3124
C: Gen. Class. 0.1339 0.2110 0.1962 0.3025 0.1028 0.0850
8 Long A: Direct Generation 0.0755 0.3220 0.3002 0.1519 0.2533 0.2757
B: Logit-Probing 0.9970 0.6293 0.5070 0.9676 0.5442 0.4164
C: Gen. Class. 0.6783 0.3903 0.3338 0.0078 0.2077 0.2478
Short A: Direct Generation 0.2636 0.1887 0.2048 0.0462 0.2108 0.1954
B: Logit-Probing 0.9652 0.3844 0.3413 0.8025 0.4415 0.2920
C: Gen. Class. 0.1429 0.2394 0.1838 0.3938 0.1122 0.1185
Table 7: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Precision. Best results for each setting are in bold. Level: L1
Qwen3-8B DeepSeek-distill-Llama-8B
Context Prediction Method Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
1 Long A: Direct Generation 0.0007 0.2977 0.3807 0.0173 0.1994 0.4070
B: Logit-Probing 0.1614 0.5224 0.7448 0.1675 0.4038 0.6063
C: Gen. Class. 0.0881 0.2823 0.5051 0.0002 0.2653 0.3910
Short A: Direct Generation 0.0089 0.2678 0.4435 0.0099 0.3070 0.4955
B: Logit-Probing 0.2604 0.5066 0.7148 0.2319 0.5153 0.6654
C: Gen. Class. 0.0304 0.2305 0.4533 0.0002 0.2313 0.3975
3 Long A: Direct Generation 0.0025 0.2850 0.4488 0.0190 0.2032 0.4138
B: Logit-Probing 0.1699 0.4997 0.7660 0.1706 0.4500 0.6296
C: Gen. Class. 0.1149 0.2773 0.5120 0.0001 0.1672 0.3889
Short A: Direct Generation 0.0180 0.2496 0.5500 0.0117 0.3146 0.5038
B: Logit-Probing 0.2899 0.5252 0.8171 0.1288 0.3325 0.5308
C: Gen. Class. 0.0304 0.2214 0.4585 0.0030 0.1171 0.4046
5 Long A: Direct Generation 0.0096 0.2624 0.4490 0.0196 0.2012 0.4378
B: Logit-Probing 0.1666 0.4903 0.7741 0.1532 0.4516 0.6763
C: Gen. Class. 0.1181 0.2950 0.5330 0.0002 0.1340 0.2853
Short A: Direct Generation 0.0926 0.2283 0.5344 0.0113 0.3070 0.5138
B: Logit-Probing 0.2859 0.5151 0.8345 0.2182 0.6206 0.7940
C: Gen. Class. 0.0272 0.2392 0.4568 0.0700 0.1174 0.1942
8 Long A: Direct Generation 0.0105 0.2411 0.4689 0.0183 0.2012 0.4208
B: Logit-Probing 0.1729 0.4966 0.7887 0.1645 0.4377 0.6477
C: Gen. Class. 0.1016 0.2811 0.4921 0.0013 0.1484 0.3565
Short A: Direct Generation 0.0605 0.2223 0.5327 0.0097 0.3088 0.5182
B: Logit-Probing 0.2870 0.5137 0.8638 0.2392 0.5892 0.7530
C: Gen. Class. 0.0301 0.2784 0.4256 0.0908 0.1279 0.2726
Table 8: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows. We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A). Each cell shows results for Long-Term (L) / Short-Term (S) preference. Metric: Recall. Best results for each setting are in bold. Level: L1

### v3

@@
-PET: Preference Evolution Tracking with LLM-Generated Explainable Distribution
+SPECTRA: Revealing the Full Spectrum of User Preferences via Distributional LLM Inference
 Abstract
-Understanding how user preference evolves over time is a fundamental challenge central to modern digital ecosystems, for which Large Language Models (LLMs) are an increasingly prominent and popular approach due to their ability to comprehend the rich semantic context within behavioral data.
-A common practice is to use LLMs to predict a user’s next action by directly generating a ranked list of preferred items.
-Although effective for short-term prediction, the end-to-end generation paradigm inherently limits personalization.
-Its opaque decision-making process obscures holistic user profiling and exacerbates popularity bias.
-To address these limitations, we propose Preference Evolution Tracking (PET), a framework that reframes the task as inferring a dynamic probability distribution over a stable and interpretable lattice of preference clusters.
-By applying logit-probing and generative classification techniques, PET infers a user’s preference as a probability distribution, enabling transparent preference learning.
-On public benchmarks (Yelp, MovieLens), PET improves ranking quality by up to $40\%$ in NDCG over direct generation baselines.
-On a large-scale, real-world dataset from a short-video platform, it excels at ranking long-tail contents, significantly outperforming a SOTA production model by $7$ times in the NDCG score.
-Ultimately, PET transforms the user profile model from direct preference list generation to a transparent distributional preference mapping, paving the way for more explainable, fair, and diverse personalization systems.
+Large Language Models (LLMs) are increasingly used to understand user preferences, typically via the direct generation of ranked item lists.
+However, this end-to-end generative paradigm inherits the bias and opacity of autoregressive decoding, over-emphasizing frequent (head) preferences and obscure long-tail ones, thereby biasing personalization toward head preferences.
+To address this, we propose SPECTRA (Semantic Preference Extraction and Clustered TRAcking), which treats the LLM as an implicit probabilistic model by probing it to infer a probability distribution over interpretable preference clusters.
+In doing so, SPECTRA reframes user modeling from sequence generation with decoding heuristics to distributional inference, yielding explicit, cluster-level user preference representations.
+We evaluate SPECTRA on MovieLens, Yelp, and a large-scale short-video platform, demonstrating significant gains across three dimensions: SPECTRA achieves (i) distributional alignment, reducing Jensen–Shannon divergence to empirical distributions by $25\%$ against strong baselines; (ii) long-tail exposure, reducing decoding-induced head concentration and increasing global exposure entropy by $30\%$; and (iii) downstream applications such as personalized ranking, translating these gains into a $40\%$ NDCG boost on public datasets and a 7× improvement on ranking long-tail preferences against an industry-leading Transformer-based production baseline.
+SPECTRA: Revealing the Full Spectrum of User Preferences via Distributional LLM Inference
+Luyang Zhang thanks: Part of this work was done while at TikTok Inc.
+Carnegie Mellon University luyangz@andrew.cmu.edu          Jialu Wang TikTok Inc.
+faldict@ucsc.edu          Shichao Zhu TikTok Inc.
+shichao.szhu@gmail.com
+Beibei Li Carnegie Mellon University beibeili@andrew.cmu.edu          Zhongcun Wang TikTok Inc.          Guangmou Pan TikTok Inc.
+Yang Song TikTok Inc. ys@sonyis.me
 1 Introduction
-Understanding user interests in a transparent and interpretable manner is a central challenge in building trusted and adaptive digital ecosystems.
-The advent of Large Language Models (LLMs) presents a transformative opportunity in this domain.
-With their profound ability to comprehend context and semantics within unstructured user behavioral histories – from product reviews to content consumption sequences – LLMs promise to move beyond simple pattern matching to a deeper, more holistic understanding of user intent.
-The ultimate goal is not merely to predict the next action of users, but to create transparent and dynamic models of user interest that can be trusted, analyzed, and fairly acted upon.
-The prevailing approach for capturing user preferences involves leveraging LLMs to directly generate ranked outputs (wang2025user; wang2024llms; ngo2024recgpt; yu2024icpl; deng2025onerec), a method proven effective for short-horizon, top-$k$ recommendation tasks.
-While effective for short-horizon prediction and identifying popular preferences, this direct-to-ranking paradigm leaves important gaps.
-First, it lacks interpretability: although LLMs encode rich internal states, these models do not expose an explicit or structured representation of the user’s overall preferences – making them difficult to audit or reuse in downstream tasks.
-Second, the focus on a small $k$, combined with popularity bias, often leads to the over-representation of mainstream content – neglecting the long-tail interests that define a user’s niche tastes and are critical for achieving holistic, unbiased personalization.
-These limitations have significant consequences for algorithmic decision-making in user-oriented applications.
-The lack of an explicit user model undermines system transparency, making it difficult to govern and audit for fairness and bias.
-Additionally, the over-reliance on popular items not only distorts model predictions but also reinforces popularity during training – suppressing the model’s ability to learn users’ nuanced, long-tail preferences.
-This feedback loop amplifies the Matthew Effect (i.e., the rich get richer), reinforcing echo chambers that reduce content diversity, entrench mainstream tastes, and marginalize underrepresented creators and user segments (bakshy2015exposure; wu2024survey).
-This motivates a paradigm shift toward transparent, distributional user modeling, leading to our central research question:
-How can we infer personalized evolving preference distribution via LLM – enhancing interpretability, fairness, and diversity?
-To address this problem, we propose Preference Evolution Tracking (PET), a framework that customizes LLMs into probabilistic inference engines for modeling users’ evolving preference distributions.
-By leveraging pre-softmax logits over a stable preference cluster lattice (e.g., categories or tags), PET captures structured relevance signals more effectively than volatile item-level spaces.
-This approach enhances interpretability and diversity, especially for long-tail preferences.
-Specifically, PET supports two scalable inference strategies: Likelihood-based Probing, which iteratively queries the model for each cluster to build a precise distribution, and the more efficient Generative Classification, which extracts all probabilities in a single forward pass.
-Furthermore, PET employs Hierarchical Probing to maintain tractable inference in extreme multi-label scenarios through semantic taxonomy traversal.
-Theoretically, the ranking according to the inferred PET distribution guarantees an optimal ranking results (e.g., with respective to NDCG/Recall) under standard assumptions.
-Applied to domain-aligned LLMs, PET produces portable, interpretable user profiles for long-horizon ranking, fairness auditing, and long-tail preference modeling.
-To validate PET’s effectiveness, we conduct comprehensive experiments across three diverse datasets with varying cluster complexity: MovieLens, Yelp, and a large-scale dataset from a world-leading short video platform.
-Our results demonstrate that the distributional approach significantly outperforms direct-generation baselines and state-of-the-art (SOTA) ranking models on public benchmarks.
-More critically, on the real-world data of the short video platform, our framework shows a unique ability to rank users’ niche, long-tail interests, significantly outperforming a SOTA model in production.
-Notably, our goal is not to replace mature production models on highly popular (head) content.
-Rather, PET complements them by recovering user-specific long-tail preferences that are typically underserved by popularity-optimized systems.
-Across all experiments, our methods prove more adept at capturing stable, long-term preferences over transient, short-term interests.
+Understanding user preferences in an explicit and interpretable manner is a central challenge in building personalized AI systems (Doshi-Velez and Kim, 2017; Amershi et al., 2019; Rudin, 2019; Lipton, 2018).
+Large Language Models (LLMs) have emerged as a promising tool due to their ability to leverage rich semantic signals from textual side information (e.g., reviews, descriptions) alongside interaction logs.
+However, moving beyond simple pattern matching benefits from treating the LLM not merely as an autoregressive generator of item candidates or ranked lists from textualized user histories, but as an implicit probabilistic model over latent user intent, in line with emerging theoretical frameworks that interpret in-context learning as approximate Bayesian inference over latent variables (Bengio et al., 2025; Wang et al., 2023; Li et al., 2023).
+This motivates seeking explicit, time-varying preference distributions over interpretable semantic clusters that can be analyzed, audited, and reused across downstream applications.
+The prevailing approach for capturing user preferences involves leveraging LLMs to directly generate ranked lists of items (Wang et al., 2025, 2024a; Ngo and Nguyen, 2024; Yu et al., 2024; Deng et al., 2025).
+Even recent methods adopting textual rationales or chain-of-thought ultimately rely on this autoregressive decoding process to produce final recommendations (Wang et al., 2024b; Yi et al., 2025).
+While effective for predicting short-term interactions with already-popular item, this paradigm leaves critical gaps.
+First, it keeps the latent preference representation implicit: generation fails to expose an explicit, structured posterior over preferences, making it infeasible for audits and explanations.
+Second, in direct generation, reliance on decoding heuristics (e.g., beam search, top-$k$ sampling) could exacerbate head concentration by truncating low-probability clusters.
+As detailed in Section˜3.3, this structural vulnerability to frequency bias and search truncation might suppresses the long tail (Holtzman et al., 2020; Stahlberg and Byrne, 2019), preventing faithful personalization of user-specific niche preferences.
+These limitations have significant consequences for user-facing AI systems.
+The lack of an explicit user model hampers transparency and governance, making it difficult to systematically inspect or audit how preferences are represented across users and groups.
+At the same time, systematically suppressing low-probability but relevant items reinforces popularity during training and deployment, reducing content diversity and marginalizing long-tail preferences on real-world platforms (Bakshy et al., 2015; Wu et al., 2024).
+Together, these observations motivate a shift from sequence generation to explicit, distributional user modeling, leading to our central research question:
+How can we infer personalized preference distributions via LLMs as explicit posteriors over semantic clusters that preserve diverse user preferences and support downstream decision-making?
+To address this question, we propose SPECTRA (Semantic Preference Extraction and Clustered TRAcking), a distributional user modeling framework that treats the LLM as an implicit probabilistic model and extracts a cluster-level probability vector over a semantically interpretable set of preference clusters by probing the model’s logits.
+Using a cluster vocabulary (e.g., tags or categories) provides a stable semantic interface, avoiding the volatility and sparsity of item-level identifiers while yielding preference representations that are comparable across users and time.
+By representing preferences as distributions – rather than decoded item lists – SPECTRA reduces head-dominated concentration and better preserves niche preferences, improving coverage and exposure diversity in the resulting preference representation.
+For scalability to large cluster space, SPECTRA remains tractable via coarse-to-fine inference over a semantic taxonomy.
+To validate SPECTRA, we evaluate on three datasets spanning varying scale and cluster complexity: MovieLens, Yelp, and a large-scale industrial short-video platform.
+Across these settings, SPECTRA consistently outperforms direct generation (DG) and strong LLM-based ranking baselines, improving distributional alignment, exposure diversity, and downstream ranking quality.
+Beyond short-horizon next-step prediction, our experiments further show that SPECTRA more reliably captures persistent, long-term preference structure over time, yielding stronger performance in long-horizon evaluation settings.
+SPECTRA achieves particularly significant gains on under-represented preference segments, consistent with its ability to mitigate head-dominated concentration; accordingly, SPECTRA is not intended to replace mature head-optimized production rankers, but to complement them by recovering user-specific niche preferences that are often suppressed under decoding-based pipelines.
 Our primary contributions are threefold:
   1. 1.
-    A New Paradigm for User Modeling: We introduce a novel framework that shifts the focus from optimizing short-term item predictions to generating interpretable, dynamic probability distributions over a stable lattice of human-understandable preference clusters.
+    Distributional user modeling from LLMs.
+We reframe LLM-based user modeling from autoregressive list generation to distributional inference, producing cluster-level preference distributions over a semantically interpretable cluster vocabulary.
   2. 2.
-    A Portable and Dynamic User Profile: The learned preference distribution serves as a rich, multi-faceted user profile that captures long-tail interests and is readily portable for diverse downstream tasks, such as user segmentation, trend analysis, or fairness audits.
+    Scalable posterior extraction via probing.
+We extract these cluster-level distributions by probing LLM logits, and design inference procedures that remain tractable for large cluster vocabularies via coarse-to-fine hierarchical inference over a semantic taxonomy.
   3. 3.
-    Theoretical and Empirical Validation: We prove that PET’s distributional ranking is theoretically superior to direct generation under standard assumptions.
-Empirically, we validate PET on three datasets (MovieLens, Yelp, short-video platform), achieving $40\%$ NDCG over direct generation and $20\%$ NDCG over a SOTA ranking model (Qwen3-Reranker-8B (qwen3embedding)) on public benchmarks, and a $7\times$ improvement in long-tail ranking against a SOTA production-level ranking baseline on a real-world dataset.
-Figure 1: The PET framework pipeline.
-Left (Training): An LLM is trained on user history to learn preference distributions.
-Center (Inference): probing methods extracts the predicted preference distribution from the model’s internal logits.
-Right (Application): The transparent distribution is used for downstream tasks: personalized ranking, long-tail discovery, and interpretable user profiling.
+    Extensive Empirical Validation.
+Across MovieLens, Yelp, and a large-scale industrial short-video platform, SPECTRA yields consistent gains in distributional fidelity and coverage: it improves alignment to empirical preference distributions (e.g., $23\%$ JS divergence vs.
+a strong LLM-based ranking baseline such as Qwen3-Reranker-8B (QwenTeam, 2025)) and increases exposure diversity (up to $30\%$ global exposure entropy) relative to direct generation.
+As a downstream consequence, these representation gains also improve ranking outcomes (up to $40\%$ NDCG on public benchmarks) and deliver up to 7× higher NDCG on under-represented preference segments than a production Transformer baseline.
+Figure 1: The SPECTRA framework pipeline.
+Left (Training): we adapt an LLM (CPT + SFT) to predict a user’s future preference distribution over $K$ clusters from historical context.
+Center (Inference): we extract cluster scores from the model’s logits via likelihood-based probing or generative classification (yielding score vector $S$, and obtain a predicted distribution $\hat{\theta}=\text{softmax}(S)$ (and ranking via $\text{argsort}(S)$).
+Application: the transparent distribution supports personalized ranking, long-tail interest discovery, and interpretable preference profiling.
 2 Related Work
-User Preference Modeling.
-To deliver effective personalization, user models aims to capture preferences that are both diverse and dynamic, through three key lines of work.
-Multi-interest models represent users via multiple embeddings linked to distinct interests (shi2023everyone; cen2020controllable; zhou2018deep).
-Sequential models, such as FPMC (rendle2010factorizing), GRU4Rec (hidasi2015session), SASRec (kang2018self), and BERT4Rec (sun2019bert4rec), focus on temporal dynamics to predict the next item in a user’s history.
-More recently, distributional modeling has become central in alignment and fairness, with methods like GDPO for group fairness (yao2024no) and RLHF-based techniques like DPL and AOT for reward uncertainty (siththaranjan2023distributional; melnyk2024distributional).
-While powerful, these alignment-based methods introduce significant practical considerations, as they often require costly preference-labeled datasets and complex, computationally intensive pipelines.
-LLMs in Preference Modeling.
-Recent work has increasingly applied Large Language Models (LLMs) to the complex task of user preference modeling.
-Early efforts leveraged LLMs as powerful components for embedding generation (U-BERT; qiu2021u) or unified the field under a prompt-based, text-to-text paradigm (P5; geng2022recommendation).
-This led to end-to-end generative models like OneRec (deng2025onerec) and (ngo2024recgpt) that directly produce ranked outputs or the most relevant items (wang2025user; wang2024llms), often incorporating sophisticated alignment modules.
-In a parallel effort toward explainability, RecGPT uses the LLM to generate descriptive, textual user profiles (yi2025recgpt).
-Crucially, the direct generation of ranked lists via LLM – the approach that forms the basis of our primary baseline – has been shown to consistently outperform traditional sequential models (like SASRec (kang2018self) and BERT4Rec (sun2019bert4rec)) in various settings (geng2022recommendation; wu2024survey; zhao2024recommender; deng2025onerec).
-Recent approaches also tackle complex scenarios through structured prompting and multi-context modeling, such as LLM4MSR (wang2024llm4msr), HUM (bao2025heterogeneous), and user interest exploration via latent clusters (wang2024llms).
-While effective for short-horizon prediction, these methods typically encode a user’s state implicitly within the model’s parameters, without surfacing an interpretable or temporally grounded preference distribution.
-Our work is also related to logit-probing and generative classification techniques for extracting model beliefs from pre-softmax logits (petroni2019language; schick2020s; raffel2020exploring; zhou2024quest), adapting them to infer evolving user preference distributions from fine-tuned LLMs.
+User Preference Representation.
+Prior work models user preferences from interaction histories using multi-preference representations (Shi et al., 2023; Cen et al., 2020; Zhou et al., 2018) and sequential next-item predictors (Rendle et al., 2010; Hidasi et al., 2015; Kang and McAuley, 2018; Sun et al., 2019).
+These methods typically encode user state implicitly (as embeddings or item-level scores), which limits transparency for analysis and auditing (Doshi-Velez and Kim, 2017; Lipton, 2018; Rudin, 2019).
+An open challenge is to surface an explicit, cluster-level preference distribution that is directly comparable across time.
+LLMs for Preference Representation.
+LLMs have been incorporated into preference representation via prompt-based unification (P5; (Geng et al., 2022)) and text-aware representation learning (U-BERT; (Qiu et al., 2021)), enabling direct list generation and generative ranking (Ngo and Nguyen, 2024; Deng et al., 2025; Wang et al., 2024a, 2025) as well as textual profile generation for explainability (Yi et al., 2025).
+Recent work enriches this with review-grounded preference cues (Kim et al., 2025), user-embedding contextualization (Ning et al., 2025), and LLM-enhanced generative retrieval (Paischer et al., 2025).
+Reasoning-style pipelines (e.g., CoT4Rec, LLMRG) generate item-level rationales for recommendations (Yue et al., 2025; Wang et al., 2024b).
+These approaches typically keep user state implicit (in hidden representations or text) or operate at the item level.
+In contrast, SPECTRA addresses the upstream task of extracting a time-comparable, cluster-level preference distribution that can serve as a structured signal for downstream rankers or reasoning modules.
+Belief Extraction and Scalable Inference.
+Belief extraction from token likelihoods has been studied mainly for factual knowledge probing and prompt-based classification (Petroni et al., 2019; Schick and Schütze, 2020; Raffel et al., 2020; Zhou et al., 2024).
+This paradigm can also be applied to infer preference distributions over semantic clusters in user modeling settings.
+Direct decoding can suffer from decoding/search artifacts (Holtzman et al., 2020; Stahlberg and Byrne, 2019), motivating inference procedures that avoid reliance on decoding heuristics.
+Distributional objectives in alignment and uncertainty modeling (Siththaranjan et al., 2023; Yao et al., 2024; Melnyk et al., 2024) further motivate predicting an explicit preference distribution.
+Scaling to large cluster spaces commonly uses coarse-to-fine hierarchical inference, as in extreme multi-label classification (Zhou et al., 2024).
 3 Theoretical and Methodological Framework
-In this section, we formalize the problem of inferring dynamic, comprehensive user preferences and introduce the Preference Evolution Tracking (PET) framework.
-We start by outlining the fundamental components of the proposed framework (see Figure 1).
-3.1 Problem Formulation
-We consider a setting in which the user’s interaction history up to time $t$ is represented as a sequence $H_{1:t}$.
-The user’s preference at any discrete time $t$ is defined as a probability distribution $\theta_{t}$ over a fixed, finite set of $K$ preference clusters $C:=\{c_{1},\cdots,c_{K}\}$:
-$\theta_{t}:=(\theta_{t}(1),\cdots,\theta_{t}(K)),\text{ with }\theta_{t}(i)\geq 0\;\forall i\in[K]\text{ and }\sum_{i\in[K]}\theta_{t}(i)=1.$ (1)
-We adopt a latent utility model (mcfadden1972conditional; luce1959individual; train2009discrete), a standard formulation in choice modeling, to describe how this true distribution arises.
-We assume that there exists a vector of unobserved “attractiveness” scores $\bm{q}:=(q_{1},\cdots,q_{K})$, where $q_{i}$ represents the true strength of the user’s preference for cluster $c_{i}$.
-The true preference distribution $\theta_{t}$ is then generated via a softmax function:
-$\theta_{t}(i)=\frac{\exp(q_{i})}{\sum_{j\in[K]}\exp(q_{j})},\forall i\in[K].$ (2)
-Our primary goal is to learn a function, parameterized by an LLM $\mathcal{M}$, that maps this interaction history to a prediction of the user’s future preference distribution, $\hat{\theta}_{t+1:T}$, for a given prediction window $\{t+1:T\}$.
-We choose to model preferences over this stable cluster space rather than the item space, as it is more robust to the high volatility and enormous scale of items in real-world digital ecosystems.
-Since this true distribution $\theta_{t+1:T}$ is not directly observable, we construct an empirical proxy for it from the user’s interactions in a time window, which we denote as $\bar{\theta}_{t+1:T}$.
-Formally, for each cluster $c_{i}\in C$, its probability is the normalized frequency of interactions:
+In this section, we formalize distributional user preference modeling in an interpretable cluster space, describe SPECTRA’s two-stage pipeline (alignment and score-based inference), and provide a brief structural analysis contrasting distributional readout with direct generation (see Figure˜1).
+3.1 Task and Setup
+Let a user’s context up to time $t$ consist of (i) an interaction history $H_{1:t}$ (e.g., rated items, reviewed businesses, watched videos), (ii) user-side metadata $U$ (e.g., bio/self-intro), and (iii) item-side metadata associated with the interacted items, denoted by $\Phi_{1:t}$ (e.g., categories/tags, prices, creator attributes).
+We therefore denote the full input context as $X_{1:t}:=\big(H_{1:t},\,U,\,\Phi_{1:t}\big).$
+We represent user preferences in a fixed, interpretable cluster space $C=\{c_{1},\dots,c_{K}\}$, where each cluster corresponds to a semantic preference category (e.g., a genre, cuisine, or topic).
+We model preferences over clusters rather than items because item spaces are large and volatile, whereas cluster semantics are stable and interpretable for tracking and auditing.
+Given context $X_{1:t}$, our goal is to predict the user’s future preference distribution over $C$ for a prediction window [t+1,T][t\!+\!1,T], denoted by $\theta_{t+1:T}\in\Delta^{K-1}$, and output $\hat{\theta}_{t+1:T}=\hat{\theta}(X_{1:t})$.
+We interpret $\theta_{t+1:T}$ as a softmax over latent utilities for clusters (equivalently, the multinomial logit form widely used in preference modeling) (McFadden, 1972; Luce and others, 1959; Train, 2009).
+Since the true $\theta$ is not directly observable, we construct an empirical target $\bar{\theta}_{t+1:T}$ by normalizing interaction counts in the future window: $\bar{\theta}_{t+1:T}(i):=n_{i}/\sum_{j}n_{j}$, where $n_{i}$ is the interaction count for cluster $c_{i}$.
+3.2 Methodology
+SPECTRA formulates user modeling as distributional inference over an interpretable cluster space.
+It consists of (1) Alignment, which trains an LLM to map user context to a preference distribution over clusters, and (2) Inference, which reads out this distribution via score-based operators, with a hierarchical strategy for large label spaces.
+(more detailed design of the methods are presented in Appendices˜A and B)
+Alignment (Continuous PT + LoRA SFT).
+We adapt the base LLM to the platform domain via continuous pre-training (CPT) on textualized domain data from user metadata, item metadata, and interaction traces $(H_{1:t},U,\Phi_{1:t})$ (e.g., serialized profiles, item titles/tags/descriptions, and interaction text such as reviews or captions).
+We then apply supervised fine-tuning (SFT) with LoRA on interaction sequences, where the target is a textual encoding of the future empirical preference distribution over clusters (computed from interactions in [t+1,T][t\!+\!1,T]), rather than a single next-item label.
+This aligns the model to predict multi-interest user preference distributions.
+Preference inference as score-based querying.
+Given context $X_{1:t}$ and a textual label/definition for each cluster $c_{i}$, SPECTRA estimates a preference distribution over the cluster space $C$ by querying the aligned LLM with pre-specified prompt templates.
+Concretely, SPECTRA constructs a score vector $S(X_{1:t})\in\mathbb{R}^{K}$ and outputs $\hat{\theta}_{t+1:T}=\mathrm{softmax}(S(X_{1:t}))$.
+We compute $S$ in two ways:
+(1) Likelihood-based probing.
+For each cluster $c_{i}$, we pose a yes/no probe (Petroni et al., 2019) under the template prompt $\pi_{\text{probe}}(X_{1:t},c_{i})$ and extract the next-token logits $\mathbf{z}_{i}\in\mathbb{R}^{|V|}$.
+Using small affirmative/negative token sets $V_{+}$ and $V_{-}$, we compute the mean logits $\mu_{i,+}:=\frac{1}{|V_{+}|}\sum_{v\in V_{+}}\mathbf{z}_{i}[v]$ and $\mu_{i,-}:=\frac{1}{|V_{-}|}\sum_{v\in V_{-}}\mathbf{z}_{i}[v]$, and set
+$S_{i}:=\mu_{i,+}-\mu_{i,-}.$ (1)
+(2) Generative classification.
+We construct a single multi-choice prompt (Raffel et al., 2020; Chung et al., 2024) $\pi_{\text{gen}}(X_{1:t},C)$ that lists clusters and maps each cluster $c_{i}$ to an answer token $v_{i}$ (e.g., A, B, …).
+The LLM produces a next-token logit vector $\mathbf{z}\in\mathbb{R}^{|V|}$ for this prompt, and we take $S_{i}:=\mathbf{z}[v_{i}]$.
+For both approaches, we eventually normalize the score vector $S$ with a softmax to obtain $\hat{\theta}_{t+1:T}$.
+Scaling via hierarchical probing.
+For extreme multi-label settings, flat likelihood probing incurs $O(K)$ scoring cost.
+We introduce hierarchical probing by organizing $C$ into a two-level taxonomy (coarse L1 categories $\rightarrow$ fine-grained L2 clusters).
+Inference proceeds top-down: (1) L1 scoping: estimate a distribution over L1 categories; (2) L2 exploration: selectively probe only L2 clusters within a subset of L1 branches (e.g., high-mass branches for relevance, or long-tail branches for exploration, depending on the application goal).
+The explored branch set $\mathcal{G}$ can be chosen to trade off relevance and exploration.
+This reduces the scoring cost to $O\!\left(K_{\mathrm{L1}}+\sum_{g\in\mathcal{G}}K_{\mathrm{L2}}(g)\right)$ (L1 categories plus the probed L2 clusters under explored branches $g$); details are in Appendix˜C.
+Hierarchical probing supports both (i) an automatic top-down expansion under high-probability parent clusters and (ii) optional analyst-guided drill-down to inspect specific preference facets; the latter is used for interpretability rather than improving ranking accuracy.
+3.3 Structural Analysis: Distributional Readout vs. Direct Generation
+We explain why SPECTRA’s distributional readout is robust relative to direct generation (DG) and summarize the key theoretical implications (proofs in Appendix˜D).
+In the idealized limit where the learned cluster scores $S(X_{1:t})$ are perfectly calibrated under cross-entropy, sorting clusters by these scores recovers the Bayes-optimal ordering for common order-aware metrics (e.g., NDCG, Recall) (Gneiting and Raftery, 2007; Goodfellow et al., 2016; Blasiok et al., 2023; Tewari and Bartlett, 2007).
+In a more realistic regime where the SPECTRA scores are only $\varepsilon$-approximately isotonic (i.e., any preferred pair is misordered with probability at most $\varepsilon$), we show that the expected number of pairwise misorderings is $O(K^{2}\varepsilon)$ (see Appendix˜D).
+Why DG degrades in practice.
+In practice, SPECTRA exhibits structural advantages over decoding-based DG.
+While both methods use the same underlying LLM, DG is susceptible to three well-documented limitations:
+  • Search / truncation.
+Decoding/search heuristics (e.g., beam search) can exhibit pathological behaviors and impose a hard truncation: valid long-tail clusters that fall outside the explored set are effectively pruned (Stahlberg and Byrne, 2019; Wiseman and Rush, 2016).
+SPECTRA instead scores clusters directly and normalizes globally, preserving tail mass.
+  • Frequency bias.
+Autoregressive generation can amplify high-frequency patterns from pre-training, biasing outputs toward head concepts (Zhao et al., 2021; Holtzman et al., 2020).
+SPECTRA normalizes scores over the given cluster set, reducing corpus-level frequency effects.
+  • Exposure bias.
+DG is sequential: early deviations change the context for later predictions, compounding errors (Ranzato et al., 2015; Bengio et al., 2015).
+SPECTRA extracts cluster scores marginally from the same user context, avoiding error compounding across positions.
+Together with the bounds in Appendix˜D, these structural considerations align with our empirical findings: as $K$ and the target list length grow, DG degrades more sharply on long-tail preference prediction, while SPECTRA remains robust.
+4 Experiment
+4.1 Experimental Design and Setup
+We evaluate SPECTRA as a cluster-level user preference modeling method: given a user history, the model infers a probability distribution over interpretable preference clusters, from which a ranked list is derived for downstream evaluation.
+We report both short-term prediction (next-step preference) and long-term prediction (future preference profile), and evaluate: (i) distributional alignment, (ii) mitigating popularity concentration collapse, and (iii) downstream ranking performance.
+Datasets.
+We use three datasets with varying cluster-space size.
+MovieLens (Harper and Konstan, 2015) offers a compact, low-dimensional space (K=19K{=}19 genres); we use 200k users with 33M ratings over 87k movies.
+Yelp represents a high-dimensional semantic space (2M users and 8M ratings across K=1,311K{=}1{,}311 business categories).
+Finally, we use a large-scale Short-Video dataset (50M interactions from 16k users over 15M items) to evaluate performance in a high-velocity, industrial setting.
+Training and evaluation protocol.
+For MovieLens and Yelp, we order each user’s interactions chronologically and split by time: the first 80% is used for CPT+SFT, and the remaining 20% is held out for evaluation.
+We report two prediction targets: long-term (LT) preference prediction, where the ground truth is the empirical cluster distribution computed over the held-out future interactions, and short-term (ST) prediction, where the target is the cluster of the next interaction.
+[^1] We define a session as a fixed-length block of consecutive interactions: 5 interactions for MovieLens, 10 for Yelp, and one day of interactions (daily aggregation) for the short-video dataset.
+For the short-video dataset, we use a fixed-window protocol: a 30-day interaction window for CPT+SFT input and the subsequent 14-day period for evaluation.
+Models and inference methods.
+We study two 8B open LMs: Qwen3-8B (QwenTeam, 2025) and DeepSeek-Distill-Llama-8B (DeepSeek-AI, 2025).
+We fine-tune each base model with our CPT+SFT pipeline using LoRA (Hu et al., 2022).
+We evaluate our primary inference methods from Section˜3.2: Likelihood-based Probing (Algorithm˜1) and Generative Classification (Algorithm˜2).
+For Yelp (large $K$), we additionally evaluate Hierarchical Probing (Algorithm˜3).
+All methods output a cluster distribution, from which we derive a ranked list over the full cluster set.
+Baselines.
+We compare against two strong LLM-based baselines (and a production personalization model for short-video).
+Direct Generation (DG) uses the same fine-tuned LLM and directly decodes a top-$k$ list of clusters (without inferring an explicit probability distribution) (Geng et al., 2022; Deng et al., 2025; Wang et al., 2025, 2024a); it is therefore evaluated on ranking and diversity metrics.
+Qwen3-Reranker-8B (Zhang et al., 2025a) is a pretrained reranker used without additional fine-tuning; it outputs per-cluster relevance scores on the same inputs, which we use for ranking and normalize into a distribution when needed.
+For the short-video dataset, we additionally compare against the platform’s production Transformer-based personalization model (a behavior-sequence model used for ranking)(Liu et al., 2022; Zhang et al., 2025b; Guan et al., 2025).
+Figure 2: MovieLens (Qwen3-8B, CPT+SFT): trends over context length (1/3/5/8 sessions) for long-term (LT; future preference profile) and short-term (ST; next-step preference) prediction.
+Left: NDCG@10 ($\uparrow$).
+Right: JS-Div to the empirical LT distribution ($\downarrow$).
+Direct Generation does not produce a distribution and is evaluated on NDCG only.
+DeepSeek-Distill-Llama-8B yields similar trends at Appendix˜F.
+Evaluation metrics.
+We report (i) distributional alignment via Jensen-Shannon divergence (JS-Div) between predicted and empirical cluster distributions, (ii) exposure diversity (mitigating popularity concentration) via Global Exposure Entropy  (Adomavicius and Kwon, 2011) computed from the cluster exposure distribution induced by users’ top-$k$ lists, and (iii) ranking performance via NDCG@k, Recall@k, and Precision@k.
+Full metric definitions and implementation details are provided in Section˜E.1.
+Computational Efficiency.
+Naive Likelihood-based Probing scales linearly with the number of clusters $K$.
+With KV-cache prefix reuse, we encode the shared user-history prefix once ($T_{\mathrm{pre}}$ tokens) and only recompute a short cluster-specific suffix per probe ($T_{\mathrm{suf}}$ tokens), which substantially reduces overhead when $T_{\mathrm{pre}}\!\gg\!T_{\mathrm{suf}}$ (see Appendix˜C).
+Our Generative Classification variant is single-shot (independent of $K$), with the same order of compute as a single forward pass.
+4.2 Robust Preference Modeling Across Context Lengths and Time Scales
+We begin with MovieLens (K=19K{=}19) to test SPECTRA across (i) input context length and (ii) prediction time scale.
+We vary the context window from 1 to 8 sessions and evaluate both short-term (next-step) and long-term (future-profile) preference prediction.
+Distributional alignment.
+Figure˜2 shows that SPECTRA yields better alignment to empirical preference distributions across context lengths, with particularly clear gains in the long-term setting.
+(DG does not output a distribution and is therefore not evaluated on JS-Div).
+For example, under CPT+SFT Qwen3-8B with context window =8=8, Logit-Probing achieves JS-Div =0.316=0.316 versus $0.406$ for Qwen3-Reranker-8B (22% relative reduction), indicating more calibrated cluster-level preference profiles.
+Pred. Method Entropy@1 Entropy@3
+LT DG 1.020 2.446
+Logit-Probing 1.747 3.130
+Gen. Class. 1.575 2.732
+ST DG 0.810 2.163
+Logit-Probing 2.093 3.246
+Gen. Class. 1.605 2.696
+Table 1: Global Exposure Entropy (higher is better) on MovieLens with CPT+SFT Qwen3-8B, reported at $k\in\{1,3\}$ for long-term (LT) and short-term (ST) prediction. Full results are in Table˜6.
+Exposure diversity and mode collapse.
+We next examine exposure diversity via Global Exposure Entropy (Table˜1).
+Across both long-term and short-term settings, the DG baseline shows pronounced concentration at small cutoffs (e.g., Entropy@1 =1.020=1.020 under Qwen3-8B, LT), suggesting decoding collapses onto a few globally popular genres.
+In contrast, SPECTRA maintains substantially higher entropy already at k=1k{=}1 (e.g., $1.747$ in LT and $2.093$ in ST), surfacing user-specific interests at the very top of the ranking.
+As expected, entropy saturates for larger $k$ due to the finite genre space, making small-$k$ behavior the most diagnostic for concentration effects.
+Downstream ranking performance.
+These profile-level improvements translate into stronger ranking, especially at larger cutoffs.
+For instance, with Qwen3-8B, context window =8=8, and long-term prediction, Logit-Probing achieves NDCG@10 =0.863=0.863, improving over DG (+55%) and Qwen3-Reranker-8B (+20%).
+While DG can be competitive at top-1, its performance degrades as $k$ increases, consistent with the concentration effects observed above.
+Our Generative Classification variant also consistently improves over DG while offering a favorable latency-accuracy trade-off via single-pass inference.
+See detailed metrics including Recall@k, and Precision@k at Appendix˜F.
+Figure 3: Group-level genre preference distributions across four time periods (MovieLens; 125 users).
+Method: SPECTRA Likelihood-based Probing with CPT+SFT Qwen3-8B.
+Darker cells indicate higher probability mass; only the top-10 genres are shown for space.
+Qualitative analysis: inspecting distributional preference shifts.
+Finally, to illustrate interpretability beyond aggregate metrics, we sample 125 users and split their histories into four equal time periods.
+Applying SPECTRA’s Likelihood-based Probing yields per-period genre distributions, whose aggregation reveals systematic changes over time (Figure˜5).
+Dominant genres such as Drama, Comedy, and Action remain relatively stable across periods, while others (e.g., Adventure, Romance, Sci-Fi) fluctuate more substantially.
+We also observe gradual increases for genres such as Sci-Fi and IMAX in the aggregated profiles, illustrating SPECTRA’s ability to surface nuanced temporal shifts in an explicit, inspectable form.
+4.3 Scalable Preference Modeling in High-Dimensional Cluster Spaces
+We next evaluate whether SPECTRA scales to high-dimensional cluster spaces using Yelp (K=1,311K{=}1{,}311 categories).
+Because flat inference over all categories is impractical at this scale, we adopt a coarse-to-fine Hierarchical Probing strategy (Algorithm˜3).
+We induce a two-level semantic taxonomy by grouping categories into $K_{1}{=}26$ L1 domains (e.g., “Food & Restaurants”, “Health & Medical”) using category text (names/descriptions); SPECTRA first infers an L1 preference distribution, then ranks L2 sub-clusters within a selected L1 branch (mapping in Appendix˜E).
+As a representative case, we report results on the high-traffic L1 domain “Food & Restaurants,” which contains $K_{2}{=}19$ L2 sub-clusters (e.g., “Mexican”, “Vegan”); other branches can be evaluated analogously.
+Notably, the taxonomy is used primarily for candidate pruning; because scoring relies on natural-language descriptors, the method the method is robust to arbitrary label IDs and to moderate changes in taxonomy design.
+Coarse (L1) identification.
+Table˜2 shows that hierarchical Logit-Probing is highly effective at identifying users’ coarse interests.
+For Qwen3-8B, it achieves NDCG@1/5/10 of 0.980/0.879/0.915 at L1, while Direct Generation degenerates (0.008/0.139/0.330), effectively failing to identify the correct coarse domain.
+This motivates coarse-to-fine inference for large $K$, and shows that the L1 branch selection step benefits most from per-cluster probing.
+Level Method NDCG
+@1 @5 @10
+L1 (26 clusters) DG .008 .139 .330
+Logit-Probing .980 .879 .915
+Gen. Class. .107 .118 .163
+L2 (19 sub-clusters) DG .298 .368 .509
+Logit-Probing .513 .562 .666
+Gen. Class. .466 .349 .454
+Table 2: Hierarchical probing on Yelp (LT, window=8=8) with Qwen3-8B.
+NDCG at L1 (26 super-clusters) and L2 (19 sub-clusters under “Food & Restaurants”).
+Best per level in bold; n=5,000n{=}5{,}000.
+DeepSeek-Distill-Llama-8B shows the same trend; see full table in Section˜F.2.
+Fine-grained (L2) ranking within a branch.
+Conditioned on the selected L1 domain, SPECTRA also improves fine-grained preference ranking among the 19 L2 sub-clusters.
+At L2, both Logit-Probing and Generative Classification outperform DG on both fine-tuned LLMs; Logit-Probing is strongest for Qwen3-8B (e.g., NDCG@10=0.666=0.666).
+DeepSeek-Distill-Llama-8B shows the same overall pattern – SPECTRA variants substantially outperform DG at both levels (see Appendix˜F).
+4.4 Long-Tail Preference Discovery in Production Short-Video Personalization
+We conclude with a large-scale, short-video dataset, comprising 78 platform-defined content clusters and a strong production Transformer-based behavior-sequence personalization model (e.g., Liu et al., 2022; Zhang et al., 2025b; Guan et al., 2025).
+This setting evaluates SPECTRA in an interaction-dense personalized system where observed engagement is shaped by an exposure feedback loop and production systems are heavily optimized for frequent head interests.We therefore focus on whether SPECTRA can complement the production stack by elevating user-specific niche interests typically under-exposed by the production stack.
+Evaluation setup.
+We aggregate user activity into daily sessions.
+Given a history window (14 or 30 days), we evaluate predictions on a held-out future window (7 or 14 days).
+Since user preference is not directly observable under strong recommendation feedback, we use cumulative play duration as a behavioral proxy for engagement and construct the ground-truth cluster profile by summing play duration over items within each cluster in the future window and normalizing across clusters (details in Section˜E.3).
+SPECTRA outputs a cluster distribution; the production system operates at the item level.
+We compare them by aggregating the clusters of served items into an exposure distribution (normalized over clusters) for the prediction window.
+Long-tail definition.
+We define the long-tail as clusters in the bottom quartile of historical engagement mass (computed on the training window), i.e., low-prevalence clusters most susceptible to under-exposure (last $25\%$ of the cluster sets, equivalent to 20 of 78 clusters in our data).
+We report NDCG@k computed specifically on this tail subset, where higher values indicate better recovery of user-specific niche interests.
+Context Prediction Method NDCG
+@1 @3 @5 @10
+14 Days 7 Days Production .007 .009 .012 .034
+Logit-Probing .113 .124 .147 .203
+14 Days Production .004 .006 .008 .024
+Logit-Probing .110 .126 .150 .208
+30 Days 7 Days Production .007 .009 .012 .034
+Logit-Probing .103 .118 .138 .192
+14 Days Production .004 .006 .008 .024
+Logit-Probing .108 .122 .143 .197
+Table 3: Short-video long-tail performance (high-activity users).
+We compare SPECTRA (Logit-Probing, Qwen3-8B with PT+SFT) against the platform production model under varying history windows (14/30 days) and prediction windows (7/14 days).
+Metric: NDCG@{1,3,5,10} on the long-tail cluster subset.
 […]
-To ensure reproducibility and simplicity, we use this transparent and normalized frequency of a single action instead of heuristic weighting of different interaction types, as our framework is designed to be agnostic to the specific construction of the ground-truth label.
-Formally, we aim to minimize the expected divergence between these two distributions:
-$\min_{\mathcal{M}}\mathbb{E}_{H_{1:t}}\left[\mathcal{D}\left(\bar{\theta}_{t+1:T}\parallel\hat{\theta}_{t+1:T}\right)\right],$ (4)
-where $\mathcal{D}$ is a measure of distributional similarity, such as the Jensen-Shannon Divergence (JS-Div).
-The ranked list of preferences used for evaluation with metrics like NDCG is then obtained by sorting the elements of $\hat{\theta}_{t+1:T}$ in descending order.
-3.2 Methodology
-This section outlines the PET framework for learning user preferences, which consists of two core stages: model alignment and preference inference.
-The alignment stage follows a structured two-phase training approach.
-First, the model undergoes large-scale pre-training (PT) on domain-specific corpora to establish foundational knowledge and capture population-level patterns.
-Subsequently, it is refined through supervised fine-tuning (SFT) on carefully curated user interaction sequences, effectively aligning its internal representations with the nuances of personalized preference prediction.
-The SFT data are constructed from each user’s full history by generating pairs (context, label), defined as follows: the context is an expanding historical window $H_{1:t}$, and the label is the user’s future preference distribution $\bar{\theta}_{t+1:T}$.
-This aligned model forms the basis for all subsequent preference inference methods.
-We now present two inference methods for preference distribution estimation, which differ primarily in their querying strategy (more detailed design of the prompts and algorithms presented in Appendices A and B).
-Method 1: Likelihood-based Probing.
-This approach decomposes a multi-class problem into a series of independent binary classifications.
-The LLM is used as a “prober” or “scorer.” For each cluster $c\in C$, we iteratively ask the LLM for inferences and extract the logit score of positive tokens and then apply a softmax function on the score vector to recover the inferred probability distribution (petroni2019language; schick2020s).
-Method 2: Generative Classification.
-This method uses a single forward pass.
-A prompt is constructed that frames the task as a multi-choice question, with each cluster mapped to a unique token (e.g., ‘A’, ‘B’).
-We then read the model’s next-token logits for these specific cluster tokens and normalize them via a softmax function to obtain the final probability distribution (raffel2020exploring; chung2024scaling).
-Extreme multi-label preference inference.
-To scale PET to real-world applications involving thousands of fine-grained preference clusters – such as in large video platforms or e-commerce systems – we introduce a two-stage Hierarchical Probing strategy.
-Existing methods face clear limitations in such settings: Likelihood-based Probing exhibits linear computational complexity in the number of clusters $K$, becoming prohibitively expensive, while Generative Classification often yields unstable and semantically ungrounded logits when forced to assign probabilities to artificial class tokens.
-To overcome these issues, we organize clusters into a two-level semantic taxonomy: coarse-grained L1 categories (e.g., Food & Drink, Technology) partition the full L2 label space into interpretable subgroups.
-This tree‑structured approach is also validated in recent extreme multi‑label classification work in other tasks (chen2023token; wan2023explore; zhou2024quest).
-Inference proceeds in two stages: (1) L1 Preference Scoping: estimates marginal probabilities over top-level categories; (2) Strategic L2 Exploration: selectively probes child clusters, guided by downstream goals.
-For example, one may prioritize top-ranked L1 branches for relevance, or intentionally explore long-tail branches for novelty.
-This decomposition improves both tractability and semantic alignment by guiding the LLM from general categories to relevant fine-grained clusters, enabling scalable and adaptive preference modeling.
-See the detailed computational complexity analysis in Appendix C.
-3.3 Theoretical Foundations and Guarantees
-This section provides a theoretical argument that PET’s methodology is optimal for standard ranking metrics compared to the direct generation paradigm. We begin with a key assumption about our model:
-Assumption 1 (Isotonic Probing Estimator).
-We assume that the vector of logit scores $S$ produced by PET’s probing methods (i.e., Algorithms 2 and 1) is order-preserving (isotonic) with respect to the true latent attractiveness scores $\bm{q}$.
-Formally, for any two clusters $c_{i},c_{j}$, we have $q_{i}\geq q_{j}\Rightarrow\mathbb{E}[S[i]]\geq\mathbb{E}[S[j]].$
-1 formalizes the goal that a successfully trained model should learn scores that are correctly ordered according to the true underlying preferences, which is a foundation for many modern preference learning models (guo2017calibration; desai2020calibration; kadavath2022language; schick2020exploiting).
-With this, we can state the following:
-Lemma 1 (Optimality of PET’s Ranking).
-With 1, the ranking produced by PET by sorting its inferred probabilities maximizes the expected score for common order-aware metrics, including NDCG@k, Recall@k, and Precision@k.
-(See proof at Appendix D.)
-Remark 1 (Sub-optimality of Direct Generation).
-The sub-optimality of direct generation stems from its autoregressive nature, where the ranking is an emergent property of a sequence generation task rather than a direct readout of preference scores.
-This process is susceptible to multiple known issues: exposure bias, where the model’s training distribution differs from its inference distribution (ranzato2015sequence); calibration errors, where output probabilities are skewed by factors like token frequency (zhao2021calibrate); and search errors from the decoding heuristics themselves, such as beam search (wiseman2016sequence).
-4 Experiment
-4.1 Experiment Design and Setup
-Our experimental design is crafted to comprehensively evaluate the efficacy of PET’s ability in discerning and predicting user preferences across varying granularities and long-term and short-term dynamics.
-We focus on using two primary inference methodologies for preference prediction and benchmark them against established and SOTA baselines.
-Datasets.
-We evaluate PET on three real-world datasets of varying scale and complexity.
-MovieLens (harper2015movielens): a widely-used dataset for movie recommendations, representing a relatively small and well-defined genre (cluster) space with 19 clusters (i.e., movie genres).
-In total, we used 200k users with 33 million rating histories associated with 87k movies.
-Yelp: a large-scale dataset with 2 million users and 8 million ratings across 1,311 business categories, providing a significantly higher-dimensional preference space.
-A Real-World Short-Video Dataset: this proprietary dataset is derived from the a real-world short video platform and captures large-scale real-world user interaction behaviors in a highly dynamic environment, reflecting diverse and rich user preference patterns.
-To ensure privacy, all identifiers are rigorously anonymized.
-The dataset comprises 50 million implicit feedback records from 16k randomly sampled users, covering 15 million items.
-Each record contains only anonymized user IDs, item IDs, and interaction types.
-[…]
-Models and Methods.
-We use two families of open-source 8B-parameter language models: Qwen3-8B (qwen3technicalreport) and DeepSeek-Distill-Llama-8B (deepseekai2025deepseekr1incentivizingreasoningcapability), selected for their strong balance of SOTA performance and computational efficiency.
-Each model is trained with our PT+SFT (LoRA fine-tune (hu2022lora)) pipeline.
-This setup supports a comprehensive ablation of training strategies.
-Our experiments evaluate the primary inference methodologies detailed in Section 3: Likelihood-based Probing (Algorithm 1) and Generative Classification (Algorithm 2).
-For the Yelp dataset with large cluster space, we specifically evaluate our Hierarchical Probing approach (Algorithm 3).
-For all methods, the final inferred probability distribution is used to generate a ranked list of clusters for evaluation.
-Baselines.
-We compare PET with two strong LLM-based baselines.
-To isolate the effect of our inference paradigm, the first baseline Direct Generation uses the same pre-trained and fine-tuned LLM as PET but is prompted to directly generate a top-k ranked list of preferences without producing a probability distribution (geng2022recommendation; deng2025onerec; wang2025user; wang2024llms).
-As this method does not produce a probability distribution, it is evaluated only on ranking metrics.
-The second baseline is Qwen3-Reranker-8B, a SOTA pre-trained model specifically optimized for ranking tasks (qwen3embedding).
-We use it out of the box without any additional fine-tuning.
-Given a user history and a candidate cluster space, it produces both a ranked list and a probability distribution.
-For the short-video dataset, we compare against the platform’s production Transformer-based ranker model.
-Evaluation Metrics.
-Model performance is comprehensively evaluated using established metrics compared to long-term (LT) and short-term (ST) ground truths, with a focus on ranking quality and distributional similarity.
-Ranking metrics include NDCG@k, which assesses the quality of ranking order with an emphasis on top positions; Recall@k, measuring the proportion of relevant items captured within the top k results; and Precision@k, evaluating the exactness of top-k recommendations.
-Distribution similarity is quantified using Jensen-Shannon Divergence (JS-Div), where lower values indicate closer alignment between predicted and actual preference distributions.
-All results are averaged across test samples and reported per context window size as specified in the dataset.
-Figure 2: Comprehensive comparison of different methods on MovieLens dataset across a range of context windows (i.e., 1,3,5, and 8 sessions).
-Prediction windows: long-term and short-term.
-Metrics: NDCG@$1,5,10$ and JS-Divergence.
-We compare our Logit Probing and Generative Classification methods against (1) Direct Generation benchmark trained (PT+SFT) on a Qwen3-8B base model and (2) a SOTA Qwen3-Reranker-8B ranking model.
-Note that for the Direct Generation baseline, the JS-divergence will not be available.
-Sample size: 38,434.
-4.2 How Effective and Robust is PET across Different Context Horizon?
-We begin by evaluating PET on the MovieLens dataset with 19 clusters, aiming to systematically test its effectiveness and robustness across following key dimensions: (1) different input history lengths and (2) prediction horizons (short-term vs.
-long-term).
-We vary the input context window from 1 to 8 sessions, and assess performance on both short-term (next-session) and long-term (multi-session) preference prediction.
-As shown in Figure 2, the results demonstrate the robustness and effectiveness of PET across a wide range of settings.
-The performance gains are especially pronounced in full-list metrics (NDCG@5, NDCG@10), where PET’s distributional inference enables richer modeling of user preferences beyond top-1 accuracy.
-For example, with Qwen3-8B, context window =8=8, and long-term prediction, Logit-Probing achieves NDCG@10 = 0.863, with a +55% improvement over Direct Generation and +20% gain over Qwen3-Reranker-8B.
-It also improves distributional alignment, reducing JS-Div to 0.316, a 22% relative reduction compared to the Qwen3-Reranker-8B’s 0.406.
-While Direct Generation is occasionally competitive at NDCG@1, its performance deteriorates as the number of predicted clusters increases, as it might suffers from exposure bias, noted in Remark 1.
-Our Generative Classification method also shows consistent improvements over Direct Generation on Qwen3-8B base model.
-Though slightly less accurate than Logit-Probing, it offers a compelling latency–accuracy trade-off due to its single-pass nature, making it a strong candidate for real-time or resource-constrained environments.
-(See more experimental results in other settings in Appendix F)
-To demonstrate PET’s ability to build interpretable profiles, we sampled 125 users and divided their histories into four equal time periods.
-Using PET’s Likelihood-based Probing, we inferred a distribution over 19 genres for each user in each period.
-Aggregating these across users yielded group-level preference distributions that evolve over time – visualized in both a heatmap and 3D ribbon plot (Figure 3).
-These results show that user preference distributions are not static but migrate in nuanced ways.
-Dominant genres like Drama, Comedy, and Action exhibit stability across periods, representing structural, long-term user interests.
-In contrast, genres such as Adventure, Romance, and Sci-Fi demonstrate significant temporal fluctuations – highlighting PET’s ability to capture volatile, context-dependent preferences.
-Notably, genres like Sci-Fi, Fantasy, and IMAX show gradual upward shifts, possibly reflecting emerging or re-surfacing interests.
-Together, these dynamics validate PET’s unique capability not only in preference prediction, but also in generating descriptive, evolving group-level user profiles – an essential feature for diagnosing user behavior shifts, designing cold-start interventions, or understanding preference stability versus exploration.
-4.3 Can PET Handle Thousands of Clusters?
-To evaluate PET’s scalability in high-dimensional settings, we use the Yelp dataset with 1,311 fine-grained categories.
-Since exhaustive Logit-Probing becomes intractable at this scale, we adopt a two-stage Hierarchical Probing strategy (Algorithm 3).
-First, we construct a semantic hierarchy that maps all categories into 26 high-level L1 clusters (e.g., “Food & Restaurants,” “Health & Medical”).
-For demonstration, we focus on the high-traffic L1 cluster “Food & Restaurants,” further decomposed into 19 L2 sub-clusters (e.g., “Mexican,” “Vegan”).
-During inference, PET first identifies a user’s coarse L1 interest, then performs fine-grained ranking within the selected L2 cluster.
-While operators may target different L1 clusters depending on downstream needs, this experiment provides one representative use case.
-(See detailed mapping construction in Appendix E.)
-Table 1 shows that PET with Hierarchical Probing achieves strong performance across both coarse (L1) and fine-grained (L2) levels.
-At L1, Logit-Probing is unequivocally dominant.
-It achieves high and robust NDCG scores (e.g., NDCG@1,5,10 are 0.980, 0.879, and 0.915 for Qwen3-8B), demonstrating an exceptional ability to identify a user’s general interests.
-In stark contrast, both the Direct Generation and Generative Classification methods fail at this stage, with performance often near zero.
-These results reinforce the necessity of iterative probing in high-dimensional settings where broad disambiguation is critical.
-At L2, where the ranking task becomes more localized, both Logit-Probing and Generative Classification significantly outperform the Direct Generation benchmark.
-Notably, while Logit-Probing remains the top performer for Qwen3-8B (0.666 NDCG@10), Generative Classification emerges as the clear winner for the DeepSeek-distill-Llama-8B model (0.574 NDCG@10), suggesting that decoding strategies for fine-grained inference can be model-dependent.
-Together, these findings demonstrate that PET, equipped with hierarchical probing, scales effectively to large cluster spaces.
-The coarse-to-fine process not only preserves tractability but also allows flexible adaptation across model architectures — with Logit-Probing excelling in breadth and Generative Classification offering a strong, efficient alternative at finer levels of granularity.
-(a) Group-level preference evolution as a probability heatmap
-(b) Group-level preference evolution as a bar chart
-Figure 3: Evolution of group-level movie genre preferences as probability distribution across four time periods.
-Dataset: Movielens, 125 users over 4 periods.
-Method: PET Likelihood-based Probing with PT+SFT on Qwen3-8B.
-Note: due to limited space, we only show top-10 clusters (genres) here.
-Qwen3-8B DeepSeek-distill-Llama-8B
-Level Method NDCG@1 NDCG@5 NDCG@10 NDCG@1 NDCG@5 NDCG@10
-L1 (26 clusters) Direct Generation .008 .139 .330 .016 .308 .343
-Logit Probing .980 .879 .915 .925 .856 .880
-Generative Classification .107 .118 .163 .001 .062 .119
-L2 (19 sub-clusters) Direct Generation .298 .368 .509 .200 .153 .329
-Logit Probing .513 .562 .666 .261 .268 .363
-Generative Classification .466 .349 .454 .684 .553 .574
-Table 1: Hierarchical Probing’s performance on the high-dimensional Yelp dataset for the long-term prediction task.
-Results are shown for a context window of 8 using the PT+SFT pipeline.
-Best NDCG scores for each level are in bold.
-L1 category contains 26 clusters; L2 category contains 19 clusters under one L1 category of “Food & Restaurants”.
-Alternative taxonomies can be easily integrated depending on platform objectives.
-Sample Size: 5000.
-4.4 How Does PET Perform on Real-World Short-Video Platform?
-We conclude our empirical study with a large-scale, private short-video dataset from a major content platform, comprising 78 content clusters.
-This experiment is designed to test our method’s efficacy in a real-world environment against a heavily optimized, production-level Transformer-based ranking model.
-The primary challenge in such an ecosystem is overcoming the inherent feedback loop, where users primarily interact with content they are already shown.
-Therefore, our evaluation focuses on PET’s ability to complement an existing production model that is heavily optimized for popular, high-frequency head content.
-In contrast, PET is designed to surface niche, long-tail interests that are often suppressed by popularity bias.
-We aggregate interactions into daily sessions (a 24-hour window), ensuring a denser and more comprehensive view of a user’s preferences.
-All results use the PT+SFT pipeline with Qwen3-8B.
-Following industry practice, we use play duration – the total time a user spends watching each content item – as a robust and continuous proxy for engagement.
-(See more details in Section E.2)
-Results in Table 2 showcase PET’s effectiveness in promoting long-tail interest, diversity, and personalized recommendation.
-We highlight the setting of 30-day history to 14-day prediction, as it represents the most challenging long-range forecasting task.
-The production model fails to identify users’ long-tail preferences, with an NDCG@20 of just 0.0243 on the long-tail segment.
-PET with Logit-Probing achieves NDCG@20 of 0.1971 on the same tail – a +711% improvement.
-This demonstrates PET’s ability to surface underrepresented, user-specific content that is typically overlooked by popularity-driven production models.
-Notably, results are consistent for $k\in\{1,5,10,20\}$ with PET achieving more gains compared to the production model when $k$ is small.
-We also compare across prediction horizons and observe that PET slightly outperforms the baseline more significantly when the forecasting window is longer (14 days vs.
-7 days), consistent with our earlier findings at Section 4.2 regarding LLMs’ strength in modeling long-term user preferences.
-Finally, PET’s performance remains robust across different context lengths.
-It maintains stable ranking quality even with limited user history, underscoring its ability to effectively summarize user interests with sparse or noisy data.
-Taken together, these findings validate PET as a practical and scalable complement to production systems.
-enhancing long-tail coverage and personal relevance without sacrificing overall performance.
-Its ability to enhance diversity and personalization – especially for the most engaged users – marks a valuable addition to real-world recommender systems.
-Context Prediction Method NDCG@1 NDCG@5 NDCG@10 NDCG@20
-14 Days 7 Days SOTA Production (Production) .007 .009 .012 .034
-Logit-Probing .113 .124 .147 .203
-14 Days SOTA Transformer (Production) .004 .006 .008 .024
-Logit-Probing .110 .126 .150 .208
-30 Days 7 Days SOTA Transformer (Production) .007 .009 .012 .034
-Logit-Probing .103 .118 .138 .192
-14 Days SOTA Transformer (Production) .004 .006 .008 .024
-Logit-Probing .108 .122 .143 .197
-Table 2: Long-tail preference learning performance on the short-video dataset for high-activity users, comparing PET’s Logit Probing against a SOTA production model across various history (i.e., 14 and 30 days) and prediction windows (i.e., 7 and 14 days).
-All results are from the PT+SFT pipeline with the Qwen3-8B model.
 5 Conclusion
-In this paper, we introduce Preference Evolution Tracking (PET), a framework that casts LLM-based personalization as distributional preference mapping: PET infers a user’s cluster-level probability vector via Likelihood-based Probing and Generative Classification, and scales to large label spaces through Hierarchical Probing.
-Under a mild isotonicity assumption, ranking by PET’s inferred probabilities is optimal for standard order-aware metrics.
-Empirically, PET is robust across datasets and horizons, outperforming strong generative and reranking baselines and substantially improving coverage of long-tail interests on an industrial-scale corpus.
-Beyond accuracy, PET yields interpretable, temporally evolving user profiles that enable group-level user analysis and preference-shift detection.
-Future work includes extending PET to model richer user representations by incorporating auxiliary signals such as demographic attributes, behavioral embeddings, and multimodal content.
-These enhancements aim to enable more personalized, adaptive, and interpretable preference modeling in real-world recommendation systems.
+We presented Preference Evolution Tracking (SPECTRA), a framework for distributional user modeling with LLMs.
+Given a user’s history, SPECTRA infers an explicit posterior over interpretable preference clusters via Likelihood-based Probing or Generative Classification, and scales to large cluster spaces with Hierarchical Probing.
+In theory, SPECTRA decouples preference inference from sequence decoding by ranking clusters using explicit probability readouts, which avoids decoding-induced exposure bias that can degrade Direct Generation at larger cutoffs.
+Across public benchmarks and an industrial short-video dataset, SPECTRA is robust across context lengths and time scales, consistently improving distributional alignment, mitigating popularity concentration, and strengthening cluster-level ranking compared to strong baselines.
+Beyond predictive gains, SPECTRA produces inspectable, temporally evolving preference profiles that support auditing and group-level analysis of preference shifts.
+Future work includes extending the framework to incorporate additional signals (e.g., multimodal content and auxiliary behavioral representations) and studying robustness under evolving taxonomies and feedback loops in deployed personalization pipelines.
 Appendix
@@
 15:  $s_{-}\leftarrow\sum_{v\in V_{-}}(\mathbf{z}_{j}[v])/|V_{-}|$ $\triangleright$ Calculate mean logit score for all negative tokens
-16:  $S[j]\leftarrow\frac{\exp(s_{+})}{\exp(s_{+})+\exp(s_{-})}$ $\triangleright$ Compute ’yes’ probability as the cluster’s score
+16:  $S[j]\leftarrow s_{+}-s_{-}$ $\triangleright$ Compute ’yes’ probability as the cluster’s score
 17:end for
@@
 […]
-7:Predicted preferences $\hat{\theta}$ over the L2 clusters $C_{L2}$.
+7:Predicted preferences $\hat{\theta}_{t+1:T}$ over the L2 clusters $C_{L2}$.
 8:
 […]
-10:Use $H_{1:t}$ and $C_{L1}$ to get L1 scores $S_{L1}$ via “Likelihood-based Probing” (Algorithm 1).
+10:Use $H_{1:t}$ and $C_{L1}$ to get L1 scores $S_{L1}$ via “Likelihood-based Probing” (Algorithm˜1).
 […]
@@
 Appendix B Prompt of Inference
-Prompt for Likelihood-based Probing Algorithm 1
+Prompt for Likelihood-based Probing Algorithm˜1
 User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
@@
 Considering the user’s long-term preferences from their movie rating history, do they like {GENRE} movies? Answer in “Yes” or “No”.
-Prompt for Generative Classification Algorithm 2
+Prompt for Generative Classification Algorithm˜2
 User History: Time 1: rated “Inception” 5/5 (Action, Sci-Fi);
@@
 Notation.
-Let $K$ be the number of preference clusters and $T$ the prompt length (tokens).
-Let $C(\mathcal{M},T)$ denote the cost of a single forward pass of the language model $\mathcal{M}$ on a prompt of length $T$ (the dominant compute).
-Each cluster $c$ is associated with a fixed “verbalizer” (token or short phrase) used to read its score from next-token logits.
-Complexity of PET Methods
-Likelihood-based Probing Algorithm 1.
-Our implementation is iterative: we issue a focused prompt per cluster, effectively scoring clusters independently. The total cost scales linearly with the number of probes:
+Let $K$ be the number of preference clusters and $T$ the total prompt length (tokens).
+Let $C(\mathcal{M},T)$ denote the cost of a single forward evaluation of the language model $\mathcal{M}$ on $T$ tokens (dominant compute).
+For KV/prefix reuse, we decompose each probing prompt into a shared prefix of length $T_{\mathrm{pre}}$ (user history + fixed scaffold) and a cluster-specific suffix of length $T_{\mathrm{suf}}$, so that $T=T_{\mathrm{pre}}+T_{\mathrm{suf}}$.
+We write $C(\mathcal{M},T_{\mathrm{pre}})$ and $C(\mathcal{M},T_{\mathrm{suf}})$ for the corresponding costs.
+Verbalizers.
+In generative classification, each cluster $c$ is associated with a verbalizer (e.g., a single token/letter) whose next-token logit is read and normalized.
+In likelihood probing, the cluster is inserted into the prompt and we read the next-token logits of a fixed response verbalizer set (e.g., Yes/No).
+Complexity of SPECTRA Methods
+Likelihood-based Probing Algorithm˜1.
+Our probing implementation is iterative: we score clusters via $K$ cluster-conditioned prompts. Without reuse, this requires $K$ full prompt evaluations:
 $\text{Cost}\;\approx\;O\!\big(K\cdot C(\mathcal{M},T)\big).$
-Generative Classification Algorithm 1.
+With prefix/KV reuse, we evaluate the shared prefix once and only re-evaluate a short cluster-specific suffix per probe:
+$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T_{\mathrm{pre}})+K\cdot C(\mathcal{M},T_{\mathrm{suf}})\big).$
+[…]
+Generative Classification Algorithm˜2.
 This method is single-shot: one prompt, one forward pass, then read the next-token logits for the $K$ verbalizers and normalize. The cost is constant with respect to $K$:
 $\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)\big).$
-Hierarchical Probing Algorithm 3.
+Hierarchical Probing Algorithm˜3.
 This two-stage method targets large $K$.
@@
 Using iterative probing at both levels, the total cost is the sum of L1 scoping and L2 exploration:
-$\text{Cost}\;\approx\;O\!\big(K_{1}\cdot C(\mathcal{M},T_{L1})\big)\;+\;O\!\big(B\cdot K_{2,\text{avg}}\cdot C(\mathcal{M},T_{L2})\big).$
+Cost $\;\approx\;O\!\big(K_{1}\cdot C(\mathcal{M},T_{L1})\big)$
+$\quad\;+\;O\!\big(B\cdot K_{2,\text{avg}}\cdot C(\mathcal{M},T_{L2})\big).$
 Since typically $K_{1}\ll K$ and $B\cdot K_{2,\text{avg}}\ll K$, this is substantially cheaper than flat iterative probing $O\!\big(K\cdot C(\mathcal{M},T)\big)$.
@@
 Direct Generation (top-$k$ list).
-One prompt with autoregressive decoding of approximately $O(k)$ output tokens:
-$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)+k\big).$
-If re-prompted separately for multiple cutoffs (e.g., $k=1,5,10$), end-to-end latency increases linearly with the number of cutoffs.
+One prompt with an initial prefill plus autoregressive decoding of $m(k)$ output tokens (typically $m(k)=O(k)$ under a constrained letter-list format):
+$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)+m(k)\cdot C_{\mathrm{dec}}(\mathcal{M})\big),$
+where $C_{\mathrm{dec}}(\mathcal{M})$ denotes the per-token decode cost given a cached prefix (implementation-dependent).
+If re-prompted separately for multiple cutoffs (e.g., $k=1,5,10$), end-to-end cost scales linearly with the number of cutoffs.
 Cross-encoder Reranker.
@@
 $\text{Cost}\;\approx\;O\!\big(K\cdot C(\mathcal{M},T)\big).$
+Practical overhead (token-length proxy).
+Define the relative compute overhead of probing (with reuse) versus a single-shot method as
+$\rho\;=\;\frac{C(\mathcal{M},T_{\mathrm{pre}})+K\,C(\mathcal{M},T_{\mathrm{suf}})}{C(\mathcal{M},T_{\mathrm{pre}}+T_{\mathrm{suf}})}.$
+In typical prompting setups, $T_{\mathrm{pre}}$ (user history) can dominate and $T_{\mathrm{suf}}$ is relatively short, so prefix reuse can reduce the overhead from $\approx K\times$ to a constant factor that depends on the ratio $T_{\mathrm{pre}}/T_{\mathrm{suf}}$.
+Absolute latency depends on batching and serving implementations; the proxy above is intended to summarize scaling rather than provide measured runtimes.
+Method # model evals No reuse With prefix/KV reuse
+Direct Generation (top-$k$) 11 prefill + $m(k)$ decode $\approx O(C(\mathcal{M},T)+m(k)\cdot C_{\mathrm{dec}}(\mathcal{M}))$ (same)
+Generative Classification 11 $O(C(\mathcal{M},T))$ (same)
+Likelihood Probing $K$ $O(K\,C(\mathcal{M},T))$ $O(C(\mathcal{M},T_{\mathrm{pre}})+K\,C(\mathcal{M},T_{\mathrm{suf}}))$
+Table 4: Complexity summary. $m(k)$ is the number of generated tokens in direct generation.
+Example regimes (token-length proxy).
+For illustration, consider K=20K{=}20 and compare probing with reuse ($T_{\mathrm{pre}}+KT_{\mathrm{suf}}$) to a single-shot pass ($T_{\mathrm{pre}}+T_{\mathrm{suf}}$) under two regimes: (i) short history: $T_{\mathrm{pre}}{=}80$, $T_{\mathrm{suf}}{=}20$, giving a ratio $\frac{80+20\cdot 20}{80+20}=\frac{480}{100}\approx 4.8\times$; (ii) long history: $T_{\mathrm{pre}}{=}300$, $T_{\mathrm{suf}}{=}15$, giving $\frac{300+20\cdot 15}{300+15}=\frac{600}{315}\approx 1.9\times$.
+These token-length ratios are proxies for scaling; actual wall-clock depends on batching and kernels.
+Regime $T_{\mathrm{pre}}$ $T_{\mathrm{suf}}$ $K$ Proxy ratio $\frac{T_{\mathrm{pre}}+KT_{\mathrm{suf}}}{T_{\mathrm{pre}}+T_{\mathrm{suf}}}$
+[…]
+[…]
+Table 5: Token-length proxy ratios for probing with prefix/KV reuse (illustrative; not measured latency).
 Practical Notes
@@
   • Multi-$k$ evaluation.
-PET’s distributional methods (Generative Classification; Hierarchical with small $B$) yield a full distribution in the same pass(es), so reporting multiple $k$ does not require re-decoding.
-Appendix D Proofs
-For Lemma 1:
+SPECTRA’s distributional methods (Generative Classification; Hierarchical with small $B$) yield a full distribution in the same pass(es), so reporting multiple $k$ does not require re-decoding.
+Appendix D Extra Theories and Proofs
+For the theoretical analysis, we consider an idealized preference mapping $g_{\phi}:H_{1:t}\rightarrow\mathbb{R}^{K}$ that produces logits $\boldsymbol{S}(H_{1:t})=g_{\phi}(H_{1:t})$ and an induced distribution $\hat{\theta}_{t+1:T}(H_{1:t})=\mathrm{softmax}(\boldsymbol{S}(H_{1:t}))$.
+We assume that $\phi$ is chosen to minimize the population cross-entropy between the latent-softmax model and the predictor:
+$\mathcal{L}_{\mathrm{CE}}(\phi)=\mathbb{E}_{H_{1:t}}\big[\mathrm{CE}\big(\theta_{t+1:T},\hat{\theta}_{t+1:T}\big)\big].$ (2)
+Training on empirical labels $\bar{\theta}_{t+1:T}$ can be viewed as a finite-sample approximation to this population objective.
+For the theoretical analysis in this appendix, we work in an idealized population setting.
+We posit a latent preference distribution $\theta(H_{1:t})=\mathrm{softmax}(\mathbf{q}(H_{1:t}))$ and view $\bar{\theta}_{t+1:T}$ as a finite-sample Monte Carlo estimate of $\theta(H_{1:t})$ (i.e., $\bar{\theta}_{t+1:T}\approx\theta(H_{1:t})$ as the number of interactions grows).
+Consequently, the empirical objective with $\bar{\theta}$ approximates the population objective with $\theta$.
+D.1 Preliminaries: Idealized Bayes-optimality
+We first establish the behavior of the model in an idealized population limit.
+Let $\boldsymbol{S}\in\mathbb{R}^{K}$ be the model’s logits, produced by SPECTRA’s probing methods (i.e., Algorithms˜2 and 1).
+The model predicts $\hat{\theta}=\mathrm{softmax}(\mathbf{S})$, and we train by minimizing cross-entropy between $\theta$ and $\hat{\theta}$.
+This follows the standard view of cross-entropy (log-loss) as a strictly proper scoring rule: at the population optimum, the predictor’s output distribution coincides with the true conditional distribution (Gneiting and Raftery, 2007; Goodfellow et al., 2016; Blasiok et al., 2023).
+In multiclass settings, this implies Bayes-consistency of the induced classifier (Tewari and Bartlett, 2007).
+Proposition 1 (Isotonicity at convergence).
+Since cross-entropy is a strictly proper scoring rule (Gneiting and Raftery, 2007; Blasiok et al., 2023), the global minimum of the population risk is achieved if and only if $\mathrm{softmax}(\boldsymbol{S})=\mathrm{softmax}(\boldsymbol{q})$.
+Equivalently, in the latent-softmax model $\theta=\mathrm{softmax}(\boldsymbol{q})$, the Bayes-optimal predictor recovers the true conditional distribution (Tewari and Bartlett, 2007; Goodfellow et al., 2016).
+By injectivity of softmax up to an additive constant, this implies $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$ for some scalar $c$.
+Thus, in the idealized limit, the model’s logits are perfectly order-preserving (isotonic) with respect to latent preferences.
 Proof.
-By 1, the scores $S$ from PET preserve the order of the true latent attractiveness scores $\bm{q}$.
-Since the softmax function is monotonic, the inferred probability distribution $\hat{\theta}$ also preserves this order.
-Thus, sorting by PET’s inferred probabilities produces a ranking, $\pi_{\text{PET}}$, that is identical to the optimal ranking, $\pi^{*}$, which sorts by the true latent scores $\bm{q}$.
-For a given user and within the time window of $\{t+1:T\}$, we let the vector of attractiveness score $\bm{q}$ be the true latent relevance score of a user for the cluster space.
-And we denote the sorted (in descending order) vector of $\bm{q}$ as $\tilde{\bm{q}}$ with $\tilde{q}_{1}\geq\cdots\geq\tilde{q}_{K}$.
-By 1, the retried logit score vector preserves the same order $S$ as $\bm{q}$.
-Thus, as the softmax function of $S$ preserves the same order of $S$, then by the procedure of PET – sorting the inferred probability distribution in the descending order – the outputted rank by the PET’s procedure has the exactly the same order as the $\tilde{\bm{q}}$.
-On the other hand, let’s visit the definition of three ranking metrics. For any given ranking/permutation of cluster set $C$, denoted as $\pi$ (with the associated score vector $q_{\pi}$), we define
-  • Recall@k: $\{\text{relevant items in top}-k\}/K$.
+Let the true latent preference distribution be $\theta=\mathrm{softmax}(\boldsymbol{q})$ and the predicted distribution be $\hat{\theta}=\mathrm{softmax}(\boldsymbol{S})$.
+The expected population cross-entropy risk is defined as:
+$\mathcal{R}=\mathbb{E}_{H}\left[-\sum_{k=1}^{K}\theta_{k}\log\hat{\theta}_{k}\right].$ (3)
+By adding and subtracting the entropy of the true distribution $H(\theta)=-\sum\theta_{k}\log\theta_{k}$, we can rewrite the risk as:
 […]
-  • NDCG@k: $\text{DCG}(\pi)@k/\text{IDCG}@k$, where $\text{DCG}(\pi)@k:=\sum_{r+1}^{K}\frac{q_{\pi}(r)}{\log_{2}(r+1)}$ and $\text{IDCG}@k$ is the best possible $\text{DCG}(\pi)@k$ over all possible permutation $\pi$.
-It is trivial to see that, all these three ranking metrics are maximized when the permutation $\pi$ has the same order as $\tilde{\bm{q}}$.
-Since the outputted rank by PET has the same order as $\tilde{\bm{q}}$, then PET achieves the optimal ranking scores with respect to these three ranking methods, which gives equal or better rank scores than the direct generation method.
+where $D_{\mathrm{KL}}$ is the Kullback-Leibler divergence.
+A fundamental property of proper scoring rules (Gibbs’ inequality) states that $D_{\mathrm{KL}}(\theta\|\hat{\theta})\geq 0$, with equality holding if and only if $\hat{\theta}=\theta$ almost everywhere.
+Therefore, the global minimum is achieved uniquely when $\mathrm{softmax}(\boldsymbol{S})=\mathrm{softmax}(\boldsymbol{q})$.
+The softmax function is invariant to constant shifts, meaning $\mathrm{softmax}(\boldsymbol{x})=\mathrm{softmax}(\boldsymbol{y})$ implies $\boldsymbol{x}=\boldsymbol{y}+c\boldsymbol{1}$ for some scalar $c\in\mathbb{R}$.
+Consequently, at the global minimum, $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$.
+Finally, since translation by a scalar $c$ does not alter the relative ordering of elements, for any pair of clusters $i,j$:
+$S_{i}>S_{j}\iff(q_{i}+c)>(q_{j}+c)\iff q_{i}>q_{j}.$ (5)
+This confirms that in the idealized limit, the logits $\boldsymbol{S}$ are perfectly isotonic with the latent utility $\boldsymbol{q}$. ∎
+In the language-modeling context, this corresponds to the standard interpretation of cross-entropy training as maximum-likelihood estimation of the next-token distribution and an upper bound on the true entropy rate (Braverman et al., 2020).
+Our analysis applies this Bayes-optimal perspective to the cluster-distribution predictor used by SPECTRA.
+D.2 Robustness: The $\varepsilon$-Approximate Regime
+Proposition proposition˜1 characterizes an ideal Bayes-optimal limit where $\boldsymbol{S}$ is exactly isotonic with $\boldsymbol{q}$.
+In practice, models are only approximately calibrated (Blasiok et al., 2023; Braverman et al., 2020).
+We capture this via a probabilistic pairwise condition.
+Assumption 1 ($\varepsilon$-Approximate Pairwise Isotonicity).
+For a given user and time window, let $\boldsymbol{q}\in\mathbb{R}^{K}$ be latent scores and $\boldsymbol{S}\in\mathbb{R}^{K}$ the logits produced by SPECTRA.
+We say $\boldsymbol{S}$ is $\varepsilon$-approximately pairwise isotonic w.r.t.
+$\boldsymbol{q}$ if, for all $i,j$ with $q_{i}>q_{j}$,
+$\mathbb{P}\Big(S_{i}<S_{j}\Big)\;\leq\;\varepsilon,$ (6)
+where the probability is over randomness in training, histories, and probing. That is, each strictly preferred pair is misordered with probability at most $\varepsilon$.
+We measure ranking quality by the number of misordered pairs relative to the Bayes-optimal ranking.
+Definition 1 (Pairwise ranking regret).
+Let $\pi^{*}$ be the permutation that sorts clusters by decreasing $q_{i}$, and let $\pi$ be any permutation of $\{1,\dots,K\}$. Define
+$\mathcal{R}(\pi;\boldsymbol{q})=\sum_{\begin{subarray}{c}i,j\in[K]\\ q_{i}>q_{j}\end{subarray}}\mathbb{I}\big(\pi^{-1}(i)>\pi^{-1}(j)\big),$ (7)
+where $\pi^{-1}(i)$ denotes the position (rank) of cluster $i$ under the permutation $\pi$ (smaller is better).
+Thus, $\mathcal{R}(\pi;\boldsymbol{q})$ counts the number of strictly preferred pairs $(i,j)$ that are reversed under $\pi$.
+Let $\pi_{\mathrm{SPECTRA}}$ be the permutation that sorts $\boldsymbol{S}$ in descending order. The following bound is immediate.
+Theorem 2 (Approximate optimality of SPECTRA).
+Under Assumption 1, the expected pairwise regret of $\pi_{\mathrm{SPECTRA}}$ satisfies
+$\mathbb{E}\big[\mathcal{R}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})\big]\;\leq\;\binom{K}{2}\,\varepsilon=O(K^{2}\varepsilon),$ (8)
+where the expectation is over the randomness in $\boldsymbol{S}$.
+Proof.
+For $q_{i}>q_{j}$, define $E_{ij}=\{\pi_{\mathrm{SPECTRA}}^{-1}(i)>\pi_{\mathrm{SPECTRA}}^{-1}(j)\}=\{S_{i}<S_{j}\}.$ By Assumption 1, $\mathbb{P}(E_{ij})\leq\varepsilon$.
+By Definition 1, $\mathcal{R}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})=\sum_{q_{i}>q_{j}}\mathbb{I}(E_{ij}).$ Taking expectations and using linearity,
+$\mathbb{E}[\mathcal{R}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})]=\sum_{q_{i}>q_{j}}\mathbb{P}(E_{ij})\leq\sum_{q_{i}>q_{j}}\varepsilon\leq\binom{K}{2}\varepsilon.$ (9)
+Thus, if SPECTRA is “mostly isotonic” in the sense of Assumption 1 (small $\varepsilon$), its ranking is close to Bayes-optimal: only $O(K^{2}\varepsilon)$ pairs are misordered in expectation.
+In the limit $\varepsilon\to 0$, we recover the exact isotonicity of Proposition 1.
+D.3 From cross-entropy to the $\varepsilon$ regime
+The previous result treats $\varepsilon$ as a behavioral property of the SPECTRA head. We now connect $\varepsilon$ to the training objective $\mathcal{L}_{\mathrm{CE}}(\phi)$.
+Let $\mathcal{L}_{\mathrm{CE}}^{\star}=\mathcal{L}_{\mathrm{CE}}(\theta_{t+1:T})$ denote the Bayes-optimal risk, achieved when $\hat{\theta}_{t+1:T}=\theta_{t+1:T}$, and define the excess cross-entropy risk
+$\Delta_{\mathrm{CE}}$ $=\mathcal{L}_{\mathrm{CE}}(\phi)-\mathcal{L}_{\mathrm{CE}}^{\star}$ (10)
+$=\mathbb{E}_{H_{1:t}}\big[\mathrm{KL}\big(\theta_{t+1:T}(H_{1:t})\,\|\,\hat{\theta}_{t+1:T}(H_{1:t})\big)\big].$ (11)
+By Pinsker’s inequality and Jensen’s inequality, we obtain:
+Lemma 1 (L1 error from excess cross-entropy).
+The expected $L_{1}$ distance between the latent and predicted preference distributions is bounded by
+$\mathbb{E}_{H_{1:t}}\big[\|\theta_{t+1:T}(H_{1:t})-\hat{\theta}_{t+1:T}(H_{1:t})\|_{1}\big]\;\leq\;$
+$\quad\sqrt{2\,\Delta_{\mathrm{CE}}}.$ (12)
+Proof.
+Recall that the excess cross-entropy risk is exactly the expected Kullback-Leibler divergence between the true distribution $\theta$ and the predicted distribution $\hat{\theta}_{t+1:T}$:
+[…]
+For any specific history $H_{1:t}$, Pinsker’s inequality bounds the $L_{1}$ distance by the KL divergence:
+$\|\theta-\hat{\theta}\|_{1}\;\leq\;\sqrt{2\,D_{\mathrm{KL}}(\theta\|\hat{\theta})}$. (14)
+Taking the expectation over histories $H_{1:t}$ on both sides:
+$\mathbb{E}\left[\|\theta-\hat{\theta}\|_{1}\right]\;\leq\;\mathbb{E}\left[\sqrt{2\,D_{\mathrm{KL}}(\theta\|\hat{\theta})}\right$]. (15)
+Since the square root function $f(x)=\sqrt{x}$ is concave, Jensen’s inequality implies that $\mathbb{E}[\sqrt{X}]\leq\sqrt{\mathbb{E}[X]}$. Applying this to the right-hand side:
+$\mathbb{E}\left[\sqrt{2\,D_{\mathrm{KL}}(\theta\|\hat{\theta})}\right$]\; $\leq\;\sqrt{2\,\mathbb{E}\left[D_{\mathrm{KL}}(\theta\|\hat{\theta})\right]}\;$ (16)
+$=\;\sqrt{2\,\Delta_{\mathrm{CE}}}.$ (17)
+Combining these steps yields the lemma. ∎
+As $\Delta_{\mathrm{CE}}\to 0$ (e.g., with more data, capacity, and optimization), the predicted distributions $\hat{\theta}_{t+1:T}$ converge to $\theta_{t+1:T}$ in expected $L_{1}$ distance.
+Intuitively, if $\hat{\theta}_{t+1:T}$ is close to $\theta_{t+1:T}$, then it is unlikely to significantly distort the ordering between clusters with a clear preference gap.
+Formalizing $\varepsilon$ as an explicit function of $\Delta_{\mathrm{CE}}$ for all pairs would require additional global margin assumptions on $\theta_{t+1:T}$, which are often unrealistic in long-tail regimes.
+Instead, we interpret $\varepsilon$ as capturing residual misorderings after training: Lemma˜1 shows that cross-entropy minimization drives $\hat{\theta}_{t+1:T}$ toward $\theta_{t+1:T}$, thereby pushing the model into a small-$\varepsilon$ regime for most histories and clearly preferred pairs.
+D.3  SPECTRA vs. Decoding-Based Direct Generation
+We now compare SPECTRA to decoding-based direct generation (DG), which uses the same logits $\boldsymbol{S}$ but produces a ranked list via an autoregressive decoding algorithm (e.g., greedy, beam, sampling).
+Let $\pi^{*}\in\Pi_{K}$ be the permutation that sorts clusters by decreasing $\boldsymbol{q}$, and let $\pi_{\mathrm{SPECTRA}}$ be the permutation that sorts $\boldsymbol{S}$.
+We model any decoding-based ranking as follows.
+Definition 2 (Decoding-based ranking).
+A decoding-based ranking procedure is a (possibly randomized) algorithm that maps logits $\boldsymbol{S}\in\mathbb{R}^{K}$ to a permutation of clusters. We denote by
+$\pi_{\mathrm{DG}}(\boldsymbol{S})\in\Pi_{K}$
+the (random) permutation produced by such a decoder when run on $\boldsymbol{S}$.
+The randomness here comes from any sampling or tie-breaking used by the decoding algorithm (e.g., top-$p$ sampling, temperature, stochastic beam search).
+We write $\mathbb{E}_{\mathrm{DG}}[\cdot]$ for expectation with respect to this decoder randomness, conditioned on the logits $\boldsymbol{S}$.
+We measure quality using a generic order-aware loss.
+Assumption 2 (Order-aware ranking loss).
+Let $\mathcal{L}:\Pi_{K}\times\mathbb{R}^{K}\to\mathbb{R}$ be a loss such that, for any fixed $\boldsymbol{q}$:
+  1. 1.
+    $\mathcal{L}(\pi^{*};\boldsymbol{q})\leq\mathcal{L}(\pi;\boldsymbol{q})$ for all $\pi\in\Pi_{K}$;
+  2. 2.
+    for almost all $\boldsymbol{q}$ (w.r.t. any continuous distribution), the minimizer is unique: $\mathcal{L}(\pi^{*};\boldsymbol{q})<\mathcal{L}(\pi;\boldsymbol{q})$ for all $\pi\neq\pi^{*}$.
+In the Bayes-optimal regime of Proposition 1, $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$, SPECTRA recovers $\pi^{*}$ by simple sorting.
+We show that no decoding procedure can do better on top of these logits.
+Theorem 3 (Dominance of SPECTRA over decoding-based generation).
+Fix $\boldsymbol{q}\in\mathbb{R}^{K}$ and assume $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$ for some $c$.
+Let $\pi_{\mathrm{SPECTRA}}=\mathrm{argsort}\boldsymbol{S}$ and let $\pi_{\mathrm{DG}}(\boldsymbol{S})\in\Pi_{K}$ be the (possibly random) permutation produced by any decoding-based procedure when run on logits $\boldsymbol{S}$.
+Under Assumption 2,
+$\mathbb{E}_{\mathrm{DG}}\big[\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})\big]\;\geq\;\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q}),$ (18)
+where the expectation is over the decoder’s internal randomness.
+Moreover, if $\mathcal{L}(\cdot;\boldsymbol{q})$ has a unique minimizer $\pi^{*}$ and the decoder outputs any $\pi\neq\pi^{*}$ with nonzero probability, then the inequality is strict.
+Proof.
+Since $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$, sorting $\boldsymbol{S}$ is equivalent to sorting $\boldsymbol{q}$: $\pi_{\mathrm{SPECTRA}}=\mathrm{argsort}\boldsymbol{S}=\mathrm{argsort}\boldsymbol{q}=\pi^{*}.$ By Assumption 2, $L(\pi^{*};\boldsymbol{q})\leq L(\pi;\boldsymbol{q})$ for all $\pi\in\Pi_{K}$.
+For any fixed realization of the decoder’s randomness, the decoder outputs some permutation $\pi_{\mathrm{DG}}(\boldsymbol{S})\in\Pi_{K}$, and thus
+$\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})\;\geq\;\mathcal{L}(\pi^{*};\boldsymbol{q})=\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q}).$ (19)
+Taking expectation over the decoder’s randomness gives
+$\mathbb{E}_{\mathrm{DG}}\big[\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})\big]\;\geq\;\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q}),$ (20)
+since $\mathcal{L}(\pi_{\mathrm{SPECTRA}};\boldsymbol{q})$ is deterministic.
+If $\mathcal{L}(\cdot;\boldsymbol{q})$ has a unique minimizer $\pi^{*}$, then whenever $\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})=\mathcal{L}(\pi^{*};\boldsymbol{q})$ we must have $\pi_{\mathrm{DG}}(\boldsymbol{S})=\pi^{*}$.
+Thus equality in expectation can hold only if the decoder returns $\pi^{*}$ almost surely.
+If instead the decoder outputs some $\pi\neq\pi^{*}$ with nonzero probability, then on that event $\mathcal{L}(\pi_{\mathrm{DG}}(\boldsymbol{S});\boldsymbol{q})>\mathcal{L}(\pi^{*};\boldsymbol{q})$, which makes the inequality strict.
 ∎
+Theorem 3 is an idealized statement: when logits are perfectly calibrated ($\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$), SPECTRA’s simple logit-sorting is Bayes-optimal, and any decoding-based mapping on top of the same logits can at best match, but not improve, the induced ranking.
+In practice, SPECTRA operates in the $\varepsilon$-approximate regime of Assumption 1, where its regret is controlled by Theorem 2, while decoding-based methods incur additional errors from exposure bias, truncation, and search heuristics.
+Remark 1 (Risk Decomposition: SPECTRA vs. Direct Generation).
+Theorem 3 characterizes an idealized regime where logits are perfectly calibrated ($\boldsymbol{S}\approx\boldsymbol{q}$).
+In the practical regime, Direct Generation (DG) incurs additional error sources that SPECTRA avoids due to its distributional nature.
+We formally conceptualize the ranking risk of DG as:
+$\mathrm{Risk}(\text{DG})$ $\;\approx\;\mathrm{Risk}(\text{SPECTRA})\;+\;\Delta_{\text{freq}}$
+$\quad\;+\;\Delta_{\text{trunc}}\;+\;\Delta_{\text{exp}}.$ (21)
+Here, $\mathrm{Risk}(\text{SPECTRA})$ is controlled by the approximate isotonicity parameter $\varepsilon$ (Theorem 2), while the additive penalty terms represent the structural limitations identified in Section 3.3:
+  1. 1.
+    $\Delta_{\text{freq}}$ (Frequency Bias): Represents the divergence caused by the autoregressive model’s tendency to over-produce high-frequency tokens from pre-training (Zhao et al., 2021; Holtzman et al., 2020), which SPECTRA mitigates via closed-set normalization.
+  2. 2.
+    $\Delta_{\text{trunc}}$ (Search/Truncation Error): Represents the expected utility loss from heuristic search (e.g., beam search), which imposes a hard truncation on valid long-tail clusters falling outside the beam (Wiseman and Rush, 2016; Stahlberg and Byrne, 2019).
+  3. 3.
+    $\Delta_{\text{exp}}$ (Exposure Bias): Represents the error accumulated due to context distribution shift during autoregressive decoding (Ranzato et al., 2015).
+This decomposition theoretically grounds our empirical observation that SPECTRA is significantly more robust in long-tail ranking tasks.
 Appendix E Detailed Experimental Design
-E.1 Yelp Dataset
-On the Yelp dataset, we evaluate PET in a high-dimensional setting with 1,311 fine-grained categories in the Yelp dataset, designed to test the scalability of our approach.
+E.1 Setups
+Global Exposure Entropy.
+Let $\mathcal{U}$ be the set of evaluation users and $\mathcal{C}$ the cluster space (e.g., genres).
+For a given method and cutoff $k$, let $R_{u,k}\subseteq\mathcal{C}$ denote the set of clusters appearing in the top-$k$ list for user $u\in\mathcal{U}$.
+We define the global exposure count for cluster $c\in\mathcal{C}$ as
+$n_{c}(k)=\sum_{u\in\mathcal{U}}\mathbb{I}\big(c\in R_{u,k}\big),$ (22)
+and the corresponding normalized exposure distribution
+$P_{\mathrm{exp},k}(c)=\frac{n_{c}(k)}{\sum_{c^{\prime}\in\mathcal{C}}n_{c^{\prime}}(k)}.$ (23)
+The Global Exposure Entropy at cutoff $k$ is then
+$\mathrm{Entropy@}k=H_{k}:=-\sum_{c\in\mathcal{C}}P_{\mathrm{exp},k}(c)\,\log_{2}P_{\mathrm{exp},k}(c).$ (24)
+Higher values of $H_{k}$ indicate that exposure is distributed more equitably across clusters, while lower values indicate concentration on a small subset of clusters (mode collapse).
+We use presence in the top-$k$ list (unweighted) to measure exposure concentration; position-weighted variants yield similar trends.
+E.2 Yelp Dataset
+On the Yelp dataset, we evaluate SPECTRA in a high-dimensional setting with 1,311 fine-grained categories in the Yelp dataset, designed to test the scalability of our approach.
 In this regime, Logit Probing becomes prohibitively expensive due to the large output space.
-To address this, we adopt Hierarchical Probing (Algorithm 3) that decomposes the task into two levels: a coarse-grained cluster identification followed by fine-grained ranking within the selected scope.
+To address this, we adopt Hierarchical Probing (Algorithm˜3) that decomposes the task into two levels: a coarse-grained cluster identification followed by fine-grained ranking within the selected scope.
 We construct a hierarchical mapping that maps 1,311 categories into 26 high-level L1 clusters (e.g., “Food & Restaurants”, “Health & Medical”).
 For a high-traffic category like “Food & Restaurants,” we then define a corresponding L2 space with 19 sub-clusters (e.g., “Mexican,” “Vegan”) under L1’s “Food & Restaurants”.
-In inference, PET first identifies the user’s broad L1 interests and then “zooms in” to perform a nuanced L2 ranking within the top-predicted L1 category.
+In inference, SPECTRA first identifies the user’s broad L1 interests and then “zooms in” to perform a nuanced L2 ranking within the top-predicted L1 category.
 While operators may freely select which high-level categories to probe based on application needs, this experiment offers one representative setup on the Yelp dataset.
@@
 With the mapping, of which we verified that Level-1 categories are well-balanced, with no single group dominating the total frequency or collapsing into overly broad labels:
-L1 Categories: Active Life & Fitness, Active Life, Sports & Recreation, Arts, Entertainment & Events, Automotive, Beauty & Spas, Cannabis Services, Community & Government, Education, Event Planning & Services, Farms & Ranches, Food & Restaurants, Health & Medical, Home & Public Services, Home Maintenance, Hotels & Travel, Internet & Communications, Local & Public Services, Nightlife & Bars, Personal Services, Pets, Professional & Financial Services, Real Estate, Religious & Community, Shopping & Retail, Specialty Shops, Specialty Vehicles L2 Categories: Bicycles, Bike Shop, Active Life, Aerial Fitness, Airsoft, Amateur Sports Teams, Archery, Badminton, Barre Classes, Baseball Fields, Basketball Courts, Beach Bars, Beaches, Bicycle Paths, Bike Parking, Bike Sharing, Bike tours, Bikes, Bocce Ball, Boot Camps, Boxing, Brazilian Jiu-jitsu, …
-E.2 Real-World Short-Video Dataset
+L1 Categories: Active Life & Fitness, Active Life, Sports & Recreation, Arts, Entertainment & Events, Automotive, Beauty & Spas, Cannabis Services, Community & Government, Education, Event Planning & Services, Farms & Ranches, Food & Restaurants, Health & Medical, Home & Public Services, Home Maintenance, Hotels & Travel, Internet & Communications, Local & Public Services, Nightlife & Bars, Personal Services, SPECTRAs, Professional & Financial Services, Real Estate, Religious & Community, Shopping & Retail, Specialty Shops, Specialty Vehicles L2 Categories: Bicycles, Bike Shop, Active Life, Aerial Fitness, Airsoft, Amateur Sports Teams, Archery, Badminton, Barre Classes, Baseball Fields, Basketball Courts, Beach Bars, Beaches, Bicycle Paths, Bike Parking, Bike Sharing, Bike tours, Bikes, Bocce Ball, Boot Camps, Boxing, Brazilian Jiu-jitsu, …
+E.3 Real-World Short-Video Dataset
 On the short-video dataset, we select the play duration as the proxy metric to represent the user interests.
@@
 Appendix F Additional Experimental Results
-F.1 Evaluation of PET on Movielens
-Inference procedure of Direct Generation. Unlike PET, which infers a full probability distribution over all 19 clusters, Direct Generation is prompted separately for each top-k values.
+F.1 Evaluation of SPECTRA on Movielens
+Figure 4: Comprehensive comparison of different methods on MovieLens dataset across a range of context windows (i.e., 1,3,5, and 8 sessions).
+Prediction windows: long-term and short-term.
+Metrics: NDCG@$1,5,10$ and JS-Divergence.
+We compare our Logit Probing and Generative Classification methods against (1) Direct Generation benchmark trained (CPT+SFT) on a Qwen3-8B base model and (2) a SOTA Qwen3-Reranker-8B ranking model.
+Note that for the Direct Generation baseline, the JS-divergence will not be available.
+Sample size: 38,434.
+Qwen3-8B DeepSeek-distill-Llama-8B
+Prediction Method Entropy@1 Entropy@3 Entropy@5 Entropy@10 Entropy@1 Entropy@3 Entropy@5 Entropy@10
+Long A: Direct Generation 1.020 2.446 3.558 3.849 0.140 1.629 2.666 3.447
+B: Logit-Probing 1.747 3.130 3.565 3.979 2.307 3.054 3.352 3.813
+C: Gen. Class. 1.575 2.732 3.103 3.705 0.952 1.926 2.608 3.329
+Short A: Direct Generation .810 2.163 3.581 3.767 0.050 1.635 2.713 3.476
+B: Logit-Probing 2.093 3.246 3.612 3.964 2.256 2.922 3.252 3.701
+C: Gen. Class. 1.605 2.696 3.211 3.681 0.073 1.972 2.350 3.433
+Table 6: Fairness performance on Movielens using the CPT+SFT pipeline.
+We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A).
+Each cell shows results for Long-Term (L) / Short-Term (S) preference.
+Metric: entropy score; context window: 3.
+Best results for each setting are in bold.
+Inference procedure of Direct Generation. Unlike SPECTRA, which infers a full probability distribution over all 19 clusters, Direct Generation is prompted separately for each top-k values.
 Qwen3-8B DeepSeek-distill-Llama-8B
@@
 Qwen3-Reranker-8B .609 .522 .638 .528
-Table 3: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows.
+Table 7: Comprehensive performance on MovieLens using the CPT+SFT pipeline across a range of context windows.
 We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A) and Qwen3-Reranker-8B.
@@
 C: Gen. Class. .907 .658 .521 .762 .597 .447
-Table 4: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows.
+Table 8: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows.
 We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A).
@@
 C: Gen. Class. .123 .422 .659 .095 .370 .547
-Table 5: Comprehensive performance on MovieLens using the PT+SFT pipeline across a range of context windows.
+Table 9: Comprehensive performance on MovieLens using the CPT+SFT pipeline across a range of context windows.
 We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A).
@@
 Best results for each setting are in bold.
-F.2 Evaluation of PET on Yelp
+(a) Group-level preference evolution as a probability 3D map
+(b) Group-level preference evolution as a bar chart
+Figure 5: Evolution of group-level movie genre preferences as probability distribution across four time periods.
+Dataset: Movielens, 125 users over 4 periods.
+Method: SPECTRA Likelihood-based Probing with CPT+SFT on Qwen3-8B.
+Note: due to limited space, we only show top-10 clusters (genres) here.
+F.2 Evaluation of SPECTRA on Yelp
+Qwen3-8B DeepSeek-distill-Llama-8B
+Level Method NDCG@1 NDCG@5 NDCG@10 NDCG@1 NDCG@5 NDCG@10
+L1 (26 clusters) Direct Generation .008 .139 .330 .016 .308 .343
+Logit Probing .980 .879 .915 .925 .856 .880
+Generative Classification .107 .118 .163 .001 .062 .119
+L2 (19 sub-clusters) Direct Generation .298 .368 .509 .200 .153 .329
+Logit Probing .513 .562 .666 .261 .268 .363
+Generative Classification .466 .349 .454 .684 .553 .574
+Table 10: Hierarchical Probing’s performance on the high-dimensional Yelp dataset for the long-term prediction task.
+Results are shown for a context window of 8 using the CPT+SFT pipeline.
+Best NDCG scores for each level are in bold.
+L1 category contains 26 clusters; L2 category contains 19 clusters under one L1 category of “Food & Restaurants”.
+Alternative taxonomies can be easily integrated depending on platform objectives.
+Sample Size: 5000.
 Qwen3-8B DeepSeek-distill-Llama-8B
@@
 C: Gen. Class. 0.0402 0.1227 0.1526 0.7341 0.1095 0.0863 0.1292 0.7328
-Table 6: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows.
+Table 11: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows.
 We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A).
@@
 C: Gen. Class. 0.1429 0.2394 0.1838 0.3938 0.1122 0.1185
-Table 7: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows.
+Table 12: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows.
 We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A).
@@
 C: Gen. Class. 0.0301 0.2784 0.4256 0.0908 0.1279 0.2726
-Table 8: Comprehensive performance on Yelp using the PT+SFT pipeline across a range of context windows.
+Table 13: Comprehensive performance on Yelp using the CPT+SFT pipeline across a range of context windows.
 We compare our Logit-Probing method (B) Generative Classification variant (C) against a Direct Generation benchmark (A).